### v3

# V2X Cooperative Perception for Autonomous Driving: Recent Advances and Challenges

Abstract
Accurate perception is essential for advancing autonomous driving and addressing safety challenges in modern transportation systems. Despite significant advancements in computer vision for object recognition, current perception methods still face difficulties in complex real-world traffic environments. Challenges such as physical occlusion and limited sensor field of view persist for individual vehicle systems. Cooperative Perception (CP) with Vehicle-to-Everything (V2X) technologies has emerged as a solution to overcome these obstacles and enhance driving automation systems. While some research has explored CP’s fundamental architecture and critical components, there remains a lack of comprehensive summaries of the latest innovations, particularly in the context of V2X communication technologies. To address this gap, this paper provides a comprehensive overview of the evolution of CP technologies, spanning from early explorations to recent developments, including advancements in V2X communication technologies. Additionally, a contemporary generic framework is also proposed to illustrate the V2X-based CP workflow, aiding in the structured understanding of CP system components. Furthermore, this paper categorizes prevailing V2X-based CP methodologies based on the critical issues they address. An extensive literature review is conducted within this taxonomy, evaluating existing datasets and simulators. Finally, open challenges and future directions in CP for autonomous driving are discussed by considering both perception and V2X communication advancements.
Index Terms:
Autonomous driving; connected and automated vehicles (CAVs); cooperative perception (CP); V2X Communication; edge computing; blockchain; 6G; digital twin; segmentation; object detection; multi-object tracking; data fusion; deep learning; artificial intelligence (AI).
I Introduction
Perception plays a pivotal role in ensuring safety in autonomous driving, facilitating awareness of the surrounding road environment and other road participants over time, and utilizing various sensor modalities [1, 2, 3, 4, 5, 6, 7, 8]. Recent advancements in intelligent vehicle technology have led to significant progress in single-vehicle perception techniques across domains, such as 3D object identification using camera [9, 10, 11, 12, 13], LiDAR [14, 15, 16, 17, 18, 19, 20], radar [21, 22, 23, 24, 25, 26], and fusion [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41], as well as 3D multi-object tracking employing LiDAR [42, 43, 44, 45, 46, 47], radar [48, 49, 50, 51], and fusion [52, 53, 54, 55, 56, 57, 58, 59], to name just a few. However, single-vehicle perception technologies still grapple with notable limitations. Sparse observations due to occlusion, restricted sensor Field of View (FoV), and sensor resolution remain primary challenges [60]. Additionally, susceptibility to sensor errors arising from adverse weather and hardware failures raises concerns about system robustness [61]. These bottlenecks constrain the potential and safety of automated vehicles in complex driving scenarios.
To address the above limitations, Cooperative Perception (CP) technology has emerged, focusing on interactions between multiple agents to overcome deficiencies in single-agent perception. The core technology of CP lies in the fusion of multi-agent sensor data, generating an accurate environmental representation, which is often seen in recent research on environmental monitoring [62], healthcare [63, 64, 65], and agriculture [66, 67, 68]. This approach has led to significant progress in the development of autonomous systems in the air [69, 70], maritime [71, 72], and on land [73, 74, 75]. CP in autonomous driving, facilitated by Vehicle-to-Everything (V2X) communications [76], offers a promising avenue to transcend the limitations of single-vehicle perception [77]. V2X communications enable Connected Autonomous Vehicles (CAVs) to enhance their perception capabilities through wireless information exchange, encompassing communication between vehicles (V2V) [78], roadside infrastructures (V2I) [79], and cloud networks (V2N) [80]. CP based on V2X significantly improves vehicle awareness on the road, addressing the challenges posed by high traffic volumes, pedestrian activity, and limited visibility. The grand vision is that V2X-based CP will become a fundamental element of future Intelligent Transportation Systems (ITSs).
(a) Without CP, each vehicle has a limited perception range. Thus, they cannot plan operations.
(b) With CP, each vehicle has a complete awareness of the traffic environment and is able to plan its operations.
Figure 1: Illustration of CP supported by V2X at an intersection.
(a) Without CP, the white hatchback has a blocked front view and thus potentially can have a fatal collision with the running boy.
(b) With CP, the white hatchback can see through the food truck, predicate the boy’s behaviour and avoid a potentially fatal collision.
Figure 2: An example of V2X CP improves road safety.
TABLE I: A Summary of the Recently Published Surveys in Cooperative Perception with V2X Technologies.
Paper Year Publication Survey Topic V2X Comm. The Taxonomy of Cooperative Perception System: Modules and Issues
EC BC DT 6G CAS PIS ACD APE CIF MH MD DRR LC DT PU TD DP S2R Sim DS
[81] 2021 IEEE Network Feature level based cooperative object detection $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$ $\surd$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
[82] 2022 IEEE IV V2I cooperative object detection and tracking $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$ $\surd$
[83] 2022 IEEE T-ITS Cooperation issues in localization, map generation, object detection and tracking $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$ $\surd$
[84] 2023 IEEE ITS-M Multi-sensor fusion and cooperative perception $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$
[85] 2023 IEEE ITS-M How to improve V2X cooperation efficiency, robustness, and safety $\times$ $\times$ $\times$ $\times$ $\times$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\surd$ $\times$ $\times$ $\surd$ $\times$ $\times$ $\times$ $\surd$
This survey 2023 -- Advancement in V2X communications supporting CP in CAVs. New taxonomy and modern generic framework of CP workflow. Challenges and future research opportunities in V2X-based CP. $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
  • Note: V2X Comm.: V2X Communication, EC: Edge Computing, BC: Blockchain, DT: Digital Twin, CAS: Cooperative Agent Selection, PIS: Perception Information Selection, ACD: Alignment for Communication Delay, APE: Alignment for Pose Errors, CIF: Cooperative Information Fusion, MH: Model Heterogeneity, MD: Modality Distinctness, DRR: Data Rate Requirements, LC: Lossy Communication, DT: Data Trustworthiness, PU: Perception Uncertainties, TD: Task Discrepancy, DP: Data Privacy, S2R: Simulation to Reality, Sim: Simulator, DS: Dataset. The exact meaning of these items in this taxonomy will be explained systematically in the following sections.
CP in autonomous driving holds immense potential, promising an enhanced user experience, elevated road safety, improved air quality, diverse transportation applications, and a multitude of cutting-edge innovations [86]. Numerous additional factors must be considered to bring V2X CP into practical application, aligning with the foundational technologies underpinning V2X CP. For example, the associated data privacy [87, 88, 89] and ethical concerns [90], the associated development of smart cities [91], urban infrastructure [92, 93], the associated standardization, regulation, and policy from governments and regulatory bodies [94, 95, 96], the associated trust issue in human-system interaction [97], and the balance between excessive reliance and information inundation [98]. Thus, while the potential of V2X CP is profound, a concerted, multi-dimensional effort is requisite for its holistic integration into our daily lives.
An example of CP supported by V2X at an intersection is illustrated in Fig. 1. It is shown from Fig. 1a that each vehicle operates individually depending on its perception systems, and no collaborations exist. Thus, each vehicle has a limited understanding of the traffic environment and can only plan its operation in a short time frame. Note that the depicted perception range in this figure is purely illustrative. In reality, the sensing range varies depending on the types of sensors and weather conditions. However, due to obstructions and varying distances, autonomous vehicles face challenges such as unreliable perception of distant objects and complete blindness or noisy observation of objects blocked from their FoV [99]. In contrast, Fig. 1b demonstrates the scenario where vehicles collaborate to achieve CP. Therefore, CP enables CAVs to understand the traffic environment holistically, eliminating blind spots and ensuring complete awareness of the traffic situation for all vehicles involved. Fig. 2 provides a specific case extracted from Fig. 1, where a white hatchback, with its front view partially obstructed by a parked food truck. As shown in Fig. 2a, the white hatchback cannot directly observe a boy chasing a football near the road. This darting-out scenario presents a significant collision risk if the white hatchback fails to stop quickly. However, when CP is active among CAVs, as shown in Fig. 2b, the white hatchback can anticipate the boy’s movement, allowing itself to slow down in advance and prepare for an emergency stop, thus preventing a potentially fatal accident. CP proves instrumental in mitigating accidents and ensuring the safety of all parties in complex traffic environments.
Numerous scholarly investigations have been dedicated to CP for autonomous driving, substantiating its benefits. However, the emergence of CP introduces new challenges in driving automation, including data trustworthiness, malicious attacks, and limited communication bandwidth. With a growing number of methodologies focused on addressing these challenges, it is imperative to conduct a comprehensive survey of the advancements in this domain, analyze remaining gaps, and provide updated research problems and directions.
Several existing surveys have reviewed the progress in V2X cooperation technologies over the past few years. These papers have reviewed previous CP approaches for autonomous driving from various perspectives. For instance, a survey on feature-level-based cooperative object detection [81] outlined research outcomes in machine learning, feature map quantification, millimetre-wave communications, and vehicular edge computing. It discussed the selection of vital feature maps to mitigate data volume and communication latency. Another work [82] reviewed the research progress of infrastructure-based CP in object detection and tracking. However, these surveys often cover only specific aspects of CP, such as applications, feature-level cooperation, or V2I collaboration, hindering readers’ systematic understanding of existing cooperation technologies. An overview of algorithms and architectures for CP was provided in [83], emphasizing critical cooperation challenges related to localization, map generation, object detection, and tracking. Additionally, the survey presented in [84] elucidated multi-sensor fusion collaborative sensing techniques from the perspectives of V2V and V2I collaborative schemes, discussing challenges in CP for autonomous driving. Despite addressing prevalent CP concerns like communication delay, location errors, information fusion, and dataset collection, these surveys lacked an examination of CP research progress concerning practical implementation challenges. Consequently, a recent work [85] assessed collaborative modules and solutions based on cooperation efficiency, robustness, and safety considerations. Additionally, these surveys have yet to cover the importance of V2X communication technology. V2X communication technology is crucial for the success of CP in autonomous vehicles and intelligent transportation systems, as it serves as the backbone architecture for transferring information between CAVs in the network, ultimately leading to CP and achieving better decision-making for CAVs on the road. Table I compares our paper with these published surveys.
This work contributes to a comprehensive understanding of CP and its enabling V2X technologies and builds upon existing survey efforts. The key contributions are as follows:
  • Filling gaps in existing surveys by conducting a thorough review of the evolution of V2X CP from its inception in 2008 to state-of-the-art proposals until the end of 2023.
  • Discussing four key enabling V2X communication technologies: edge computing, blockchain, digital twins, and 6G communications.
  • Introducing a new taxonomy to review V2X CP solutions, addressing significant practical driving scenario issues.
  • Presenting a unified framework for V2X CP to elucidate essential components of multi-agent collaboration, enhancing researchers’ systematic understanding of the CP system.
  • Summarizing lessons learned, providing insights into existing challenges, and highlighting potential avenues for future research.
Figure 3: Organization of this paper.
The structure of this paper is depicted in Fig. 3. Section II provides a brief historical overview of CP system development. Section III discusses recent advances in V2X communications, the foundational infrastructure for CP. Section IV presents a generic framework for comprehending the CP system systematically. Technical issues related to CP systems and associated datasets are summarized in Sections V and VI, respectively. Section VII explores open challenges and potential future directions, followed by the conclusion in Section VIII. A list of key acronyms and abbreviations used throughout the paper is given in Table II.
TABLE II: List of Key Acronyms.
Acronyms Definitions
AI Artificial Intelligence
BEV Bird’s Eye View
CAM Cooperative Awareness Message
CAV Connected Autonomous Vehicle
CDA Cooperative Driving Automation
CI Covariance Intersection
CNN Convolutional Neural Network
CP Cooperative or Collaborative Perception
CPM Collective Perception Message
DT Digital Twin
FL Federated Learning
FoV Field of View
GNN Graph Neural Network
GPS Global Positioning System
IMU Inertial Measurement Unit
ISAC Integrated Sensing and Communication
ITS Intelligent Transportation System
LSTM Long Short-term Memory
MIMO Multiple-input Multiple-output
RSU Roadside Unit
V2I Vehicle-to-Infrastructure
V2V Vehicle-to-Vehicle
V2X Vehicle-to-Everything
II Development of Cooperative Perception: From Early Time Exploration till Recent Research
CP involves the exchange of detected obstacles or perception data among agents, thereby expanding their perceptual range, enhancing situational awareness, and ensuring safety [100]. In the context of driving automation, Cooperative Prediction entails receiving intention or desired trajectory information from neighbouring vehicles, thereby enhancing the accuracy and efficiency of motion prediction. This approach offers the potential to improve strategic ego-planning functionality [101].
Figure 4: Overview of collaboration schemes for autonomous driving.
In its nascent phase, research in cooperative transportation systems primarily centred on devising statistical signal processing and rule-based algorithms, often supplemented by data from artificial sources or other sensors such as radar and LiDAR. For example, a study outlined in [102] introduced a particle filter-based multi-target tracking framework that expanded detection range by utilizing artificial radar measurements and facilitated vehicle position sharing via V2V communication. Simulation outcomes underscored CP’s potential to enhance the performance of intelligent vehicle systems by amalgamating information from multiple vehicles, thus complementing single-vehicle detection systems. Similarly, in [103], research aimed to construct a synthesized representation of the surrounding environment for multi-vehicle CP. This was achieved by transforming the visual perspectives of two vehicles into a shared spatial reference system predicated on the estimated relative positions between the two vehicles. Subsequently, a laser line was employed to generate an approximate depth map of visual perceptions, establishing correspondence between the images. This innovative algorithm conferred the ego vehicle with the ability to perceive objects obstructed in the foreground. Nevertheless, it had notable limitations, primarily stemming from imprecise pose estimation and the use of an approximate depth map.
CP can be viewed as a mechanism designed to tackle the challenge of map merging, aligning perception data across multiple agents, and integrating it into a consistent reference frame [104]. Achieving precise results in this context necessitates temporal and spatial alignment before the post-data fusion phase. Early studies addressed this challenge by employing motion models and estimation algorithms rooted in statistical signal processing. In [104], researchers leveraged a Car2X-based perception module for object-level data fusion, utilizing an Unscented Kalman Filter to align the received bounding box proposals from communicating vehicles and roadside stations. This module facilitated the prediction of the object’s state directly to the current time step, utilizing the Constant Turn Rate and Acceleration motion model within the spatial coordinate frame of the objects. Similarly, in the realm of cooperative tracking for multiple vehicles, the study presented in [105] extended a Gaussian Mixture Probability Hypothesis Density (GM-PHD) filter with a collaborative fusion algorithm. State propagation relied on the Constant Turn Rate and Velocity motion model, accommodating variable delays in wireless communication by condensing the communicating tracks into GM-PHD densities. It is noteworthy, however, that these motion models, while effective in their own right, impose significant constraints that limit their applicability within the domain of vehicular motion.
The rule-based algorithm, devoid of a learning process, offers a computationally efficient approach with demonstrated engineering viability. In [106], an engineering-focused multi-modal CP system was introduced to facilitate cooperative autonomous driving. This system employed a mirror-neuron-inspired intention awareness algorithm. The mirror neuron concept refers to neurons that activate when an organism performs or observes a similar action in others, forming the basis for generating intention awareness. Specifically, it enabled all vehicles’ collective synthesis of anticipatory trajectories by simulating drivers’ behaviours. However, it is essential to note that this approach relied heavily on precise estimation outcomes from each agent within the vehicular network due to its use of track-level data fusion, potentially affecting model effectiveness.
Many conventional methods face challenges in effectively extracting attributes from shared information that accurately represent the necessary entities for fulfilling the requirements of driver assistance systems. In recent years, deep learning technology has demonstrated significant advancements in single-vehicle perception. As a result, numerous researchers have been actively involved in developing learning-based CP frameworks to overcome the limitations of traditional approaches. For instance, in the realm of cooperative object detection and classification, researchers have leveraged advanced deep learning techniques such as the YOLO network [107] and DenseNet [108] to extract 3D information from shared objects [109]. Further discussion on recent research in this area will be presented in Sections IV and V.
Contemporary CP strategies encompass three collaboration styles: early, intermediate, and late collaboration, each characterized by the nature of shared data. To enhance the efficacy of multi-agent cooperation, certain studies have embraced hybrid collaboration approaches, as exemplified in [110], combining early and late collaboration to consolidate information from diverse infrastructures. Fig. 4 provides a schematic overview of these collaboration schemes, considering a scenario involving two CAVs.
As illustrated in Fig. 4, early collaboration involves aggregating raw sensor observations collected by all participating agents, fostering a comprehensive perspective [111, 112, 113]. This approach is capable of addressing challenges like occlusion and long-range obstacles encountered in single-agent perception. However, its practical implementation necessitates substantial communication bandwidth. In contrast, late collaboration involves exchanging perceptual outcomes, requiring minimal data transmission [114]. While efficient in communications resource usage, late collaboration may introduce inherent noise and incompleteness into individual perceptual outputs, potentially resulting in suboptimal fusion outcomes. In response to the trade-off between performance and bandwidth constraints, intermediate collaboration has emerged as a viable solution. This approach aggregates middle features across multiple agents, incorporating representative information within compressed features. This integration holds promise for enhancing both communication efficiency and perception capabilities.
Traditional approaches in CP have typically relied on late cooperation strategies due to the complexity of feature extraction and limited communication capabilities. These strategies operate at the object level or track level. However, with the rapid progress of deep learning, intermediate collaboration has gained prominence in multi-agent cooperation, primarily due to its ability to alleviate the burden on communication resources [115]. Nevertheless, it risks introducing extraneous and unnecessary information, particularly in dynamic networking conditions. To address this challenge, optimizing cooperative information and collaborator selection becomes imperative within a CP system. Furthermore, deep representation introduces higher levels of abstraction and implicitness in the spatial interactions between agents, necessitating suitable data alignment and information fusion algorithms. Developing a unified CP framework to address these issues is of significant importance in advancing Cooperative Driving Automation (CDA).
III V2X Communication
V2X communication is the fundamental architecture that enables CP applications for autonomous vehicles. This section reviews recent advancements in V2X communications for CP applications, focusing on four enabling technologies: edge computing, blockchain, digital twins, and 6G communications.
III-A Edge Computing-enabled V2X Communications
In modern CP systems, addressing the computational and storage requirements of extensive perception data gathered from vehicular sensors and devices is essential. Cloud computing and data centres have emerged as prevalent solutions, utilizing V2X communications to interconnect vehicular devices and servers for data sharing. However, these solutions can introduce significant communication overhead and offer limited computational flexibility for real-time CP tasks. A more efficient alternative is edge computing, which positions data and computation services at the network’s edge, in close proximity to vehicular devices, facilitating rapid response in CP operations [116].
As an example, edge computing has been proposed for V2X communication links in [117] to address computation-intensive tasks and facilitate timely decisions for data offloading, a critical requirement in real-time CP scenarios due to its proximity to end users. A low-latency task offloading model was established, allowing for the offloading of CP multi-tasks via V2X uplinks, managed by a mobile edge computing server positioned at Roadside Units (RSUs). A mobility-aware dynamic offloading strategy was introduced to account for the dynamic nature of vehicles during task offloading, capturing time-varying characteristics such as vehicular mobile positions, V2X communication status, and data task specifics.
In a similar work, computation offloading was employed to enable rapid-response CP services by integrating V2X communications into edge computing [118]. Specifically, a novel joint optimization framework was presented, addressing the allocation of caching space and computation resources for vehicular edge servers. The aim was to minimize offloading latency by considering factors such as offloading decisions and edge computing capabilities, including maximum bandwidth resources and available data storage. Beyond data offloading, edge computing also provides a viable solution for low-latency content caching, with edge servers, including base stations and RSUs, employed to collect and process data offloaded by nearby vehicular devices in real-time [119]. In this context, efficient resource management, encompassing resource allocation for V2X communication channels and load balancing among edge servers, is pivotal in ensuring effective content caching for CP applications.
Clustering presents an intriguing approach to further mitigate communication latency in V2X links, where edge computing services are partitioned into distinct zones, with each zone exclusively serving a specific number of connected vehicles to enhance CP task efficiency [120]. In this setup, each CP task can be managed by an RSU master within its coverage area, while computing tasks from vehicles in other zones are routed to the appropriate RSUs to prevent communication interference. In a related study, efforts were made to minimize V2X communication latency in vehicular services, including CP, by jointly considering both 5G-enabled uplink and downlink delays, data retrieval overheads, and content execution latency [121]. Latency assessments were conducted under diverse vehicular 5G communication configurations, demonstrating the advantages of edge computing in terms of content execution time savings compared to alternatives like cloud computing.
Another investigation on latency for edge computing in V2X communication was introduced in [122], where a radio latency model was formulated to depict the interrelation between backhaul latency at the edge server and data transmission latency between the edge server and end users based on Signal-to-Noise Ratio measurements at the evolved NodeB. Simulation results unveiled the positive correlation of vehicle density, edge server location, and data execution delays, providing valuable insights into integrating edge computing services into V2X communications to minimize data content processing time.
Moving from a single-edge computing scenario, the work in [123] extended the concept to a multi-edge computing framework, where distributed edge servers collaboratively serve the shared network of end users in a three-layer V2X architecture. At the lowest layer, vehicles and end users can offload their data to execute CP algorithms hosted at the edge server, managed at the second layer. Given the intricate nature of edge computing resources, an energy-aware edge service selection algorithm was developed to choose the suitable edge server for service requests from end users. In extensive data processing cases, end users would offload to the cloud server at the third layer. To effectively handle various data requests, a task priority classification scheme was integrated to ensure equitable data execution among different end-user groups.
III-B Blockchain-enabled V2X Communications
In addition to latency considerations, security is paramount in information exchange and data sharing over V2X communications among CAVs and agents within CP systems. Inadequate data protection and security measures can leave communications vulnerable to attacks, potentially leading to information theft and posing critical challenges to the viability of CP applications [124]. Another vulnerability arises from the traditional CP networking topology, where a centralized server predominantly manages perception algorithms. A breach of this central server can result in the collapse of the entire CP system. Consequently, there is a pressing need for a more decentralized CP solution, with blockchain technology emerging as a promising candidate due to its secure and decentralized data ledger capabilities.
Blockchain offers unique security attributes, such as decentralization, availability, traceability, and trustworthiness, presenting intriguing solutions encompassing decentralized V2X communications, distributed collaborative information perception among CAVs, and information verification and user authentication for secure CP services [125]. For instance, a proof-of-concept was presented in [126], utilizing blockchain for decentralized V2X communications. Vehicular data and messages from CAVs are encrypted and stored as transactions within the blockchain, distinguishable by vehicle identification and public keys. Rather than relying on a central authority, blockchain facilitates the creation of a message management ecosystem where distributed vehicles possess equal rights to monitor and verify information through blockchain accounts embedded within their software. Smart contracts, self-executing software running atop the blockchain, were employed to establish a straightforward mechanism for decentralized governance and information authentication [127].
In [128], a Proof-of-Local Learning mechanism was introduced, encrypting and sharing data collected by CAVs through a chain of blocks monitored by a distributed network of RSUs. Blockchain facilitates equitable information exchange and low-latency CP information storage via edge computing. Blockchain technology was also harnessed in [129] to enhance secure V2X communications within a three-layer CP system, encompassing the perception, network, and blockchain layers. The perception layer includes an Android auto application for CAVs, offering parking space and vehicle localization services. The network layer facilitates interconnection between vehicles, RSUs, and the cloud server, while the blockchain layer delivers security services, including immutability, integrity, and confidentiality for information exchanged across the other two layers. Smart contracts were integrated to provide traceable payment for vehicles, offering a mobile transaction cost calculation mechanism that ensures equitable payment for data storage and retrieval requests. Furthermore, blockchain-based smart contracts were employed in [130] for V2X links, showcasing their capability to address jamming attacks aimed at intercepting messages shared over V2X links, even when encrypted. The tracing capability of blockchain technology enables tracking shared information, including its source and destination, through transaction logs. NS-3 simulator simulations showed that blockchain allows for low-latency information exchange and V2X communication in a decentralized networking environment.
Moreover, in a study by [131], a blockchain-based transmission solution for V2X communications was introduced, focusing on the block mining mechanism. Blockchain mining is essential for data verification but can be energy-intensive and result in significant communication overhead. To address this, an offloading methodology was proposed, enabling the offloading of mining tasks to mining clusters rich in computing resources dedicated to transaction verification. Meanwhile, perception messages among CAVs can be exchanged without incurring additional latency costs, aiming to deliver a high-quality user experience in CP applications. In addition to data protection, blockchain can offer intrusion detection capabilities to enhance the security of V2X communications [132]. A micro-blockchain scheme was devised, facilitating reliable intrusion detection through a joint design involving one macro-blockchain and several micro-blockchains. The quality of local intrusion samples, a key factor for efficient intrusion detection, is securely managed through a tamper-resistant ledger integrated into the vehicle network. However, when deploying blockchain for secure V2X communications in CP applications, addressing various critical technical challenges, including scalability, high transaction latency, and limited flexibility [133] stemming from complex ledger settings, is crucial. Comprehensive solutions need to be further explored to ensure robust blockchain-based V2X communications.
III-C Digital Twin-enabled V2X Communications
In V2X communications for CP, a large amount of data is collected from surrounding environments, including vehicles and RSUs. To enhance the efficiency of CP algorithms, digital reproduction of the physical V2X environment by involving such data would be an attractive solution where Digital Twin (DT) emerges as an ideal technology candidate. DT-empowered V2X networks allow the creation of a cyber 3D platform to represent a physical V2X system by combining multi-modal sensory data such as messages and signals from CAVs and CP agents.
For instance, a DT-based V2X system introduced in [134] integrated a 3D virtual urban environment and traffic simulation into an information perception system comprising distributed vehicles and a base station. To construct an urban mobility simulator, this system gathered perceptive traffic information, encompassing various elements such as stochastic vehicle types, traffic light systems, and automated driving behaviours. This DT-based solution reduces testing costs associated with on-road experiments. It provides a practical environment for evaluating the quality of V2X communications, including channel estimation, which can be implemented using neural networks. In a different context, a DT-based V2X communication system founded on vehicle-to-cloud computing was tailored for a CP-enabled cooperative ramp merging application [135]. Field implementation was conducted with three typical passenger vehicles, demonstrating that the proposed system enhances current ramp merging scenarios’ safety and environmental sustainability. A realistic DT model was constructed to manage computations at the cloud server, coordinating data sharing from distributed vehicles via 4G networks. A CP simulation system was established within this framework, enabling virtual drivers to control their vehicles with DT support, facilitating safe and smooth cooperative merging maneuvers within the same coverage area.
DT technology also found application in V2X communication visualization through multi-vehicle experiments [136], incorporating both physical and virtual vehicles to execute collaborative perception tasks. A DT model was introduced, featuring two designs: cyberspace with twin objects and physical space with a localization testbed. Cameras captured vehicle motion images, which were shared with the cloud server via V2X communications for motion analytics, including the position and heading of vehicles. Furthermore, DT plays a role in assessing the reliability of V2X communications. In [137], a site-specific DT framework was introduced to evaluate the error rate of vehicular wireless links through a combination of geometry-based channel model simulation and realistic hardware-in-the-loop experiments. Reflective objects such as traffic signs and vehicles were employed to simulate data transmission in both simulations and testbeds, demonstrating a strong correlation in transmission error rate estimation.
To cope with multi-modal V2X communications where vehicular data such as traffic signals and vehicle messages are multi-modal, a DT-based framework was introduced in [138] as high-definition 3D mapping for physical V2X communication environments. Ego vehicles are equipped with multiple sensors, such as cameras, inertial measurement units, and radars, to perceive information from surroundings. The DT model was developed through three stages: sensory data acquisition from sensors, semantic instance segmentation, and digital 3D reconstruction. Several evaluations were conducted on an unreal engine-based automotive simulator for the DT model, showing feasible V2X channel modelling and efficient sensory data visualization. This research is essential in providing further insights into using DT for high-quality V2X communication analytics and perceptive information processing. Another similar line of research is reference [139], which suggested a DT-based CAV communication model by considering V2X communications and driving behaviour patterns. This framework was useful for forming cooperation of distributed drivers, especially at non-signalized intersections, where a cooperative motion control solution was integrated to generate the vehicle speed trajectories to achieve traffic flow consensus concerning communication constraints, including delay and packet loss.
TABLE III: Summary of the Works of V2X Communications.
Category Method Year Scenario Contribution
Edge Computing [117] 2021 V2X An edge computing model was proposed for low-latency task offloading in V2X communications. A mobility-aware dynamic offloading strategy was introduced to handle the dynamics of CAVs during task offloading.
[118] 2019 V2V A computation offloading was proposed to enable fast-response CP services by integrating V2X communications into edge computing.
[121] 2019 V2X An optimization framework was developed to minimize V2X communication latency, including CP, in vehicular services by jointly considering 5G-enabled uplink and downlink delays, data retrieval overheads, and content execution latency.
[123] 2021 V2V A multiple-edge computing framework was presented in a V2X communication setting, where distributed edge servers are jointly employed to serve the shared network of end users in a three-layer V2X architecture.
Blockchain [126] 2021 V2X A proof-of-concept was proposed using blockchain for decentralized V2X communications. The vehicular data and messages from CAVs can be encrypted and stored in the blockchain as a ledger of transactions, which are distinguished by vehicle identification and public keys.
[128] 2022 V2V A proof-of-local learning mechanism was proposed where data collected by CAVs are encrypted and shared by a chain of blocks jointly monitored by a distributed network of RSUs.
[130] 2023 V2X A blockchain-based smart contract scheme was provided for V2X links. Blockchain was used to cope with jamming attacks, aiming to steal messages shared over V2X links via an immutable data ledger.
[132] 2019 V2V A micro-blockchain scheme was built to offer reliable intrusion detection enabled by a joint design of one macro-blockchain and several micro-blockchains.
Digital Twin [134] 2022 V2X A DT-based V2X system was proposed to integrate the 3D virtual urban environment and traffic simulation in an information perception system consisting of distributed vehicles and a base station.
[135] 2022 V2V A DT-based V2X communication system was considered using vehicle-to-cloud computing for a CP-enabled cooperative ramp merging application.
[136] 2022 V2X DT was investigated in V2X communication visualization through multi-vehicle experiments, using both physical and virtual vehicles to execute collaborative perception tasks.
[137] 2022 V2X A DT-based CAV communication model was proposed by taking into account V2X communications and driving behaviour patterns.
6G [140] 2022 V2X A trial using 6G for V2X communications was introduced where intelligent reflecting surfaces were integrated to facilitate V2X communications by addressing signal blockage and large-scale fading due to other vehicles and ubiquitous objects in urban cities.
[141] 2023 V2X 6G-based intelligent reflecting surfaces were exploited to optimize the achievable transmission rate for V2X communication links between the base station and CAVs.
[142] 2023 V2X A resource allocation strategy for 6G-based V2X communications was studied based on reinforcement learning, which enabled the execution of allocation decisions in an environment created by the interaction of the base station and end users.
[143] 2022 V2X A cooperative collision detection mechanism was proposed, where reinforcement learning was applied to establish a near-optimal policy for vehicle trajectories and V2X communications in a 6G wireless network setting.
III-D 6G-enabled V2X Communications
Enabled by the revolution of wireless communication technology with 6G as an emerging wireless generation, V2X communications in CP applications are envisioned to be significantly upgraded with many new interesting features offered by 6G [86]. These would be ultra-low communication latency, high data rate, and connection reliability, with improved energy and spectrum efficiency compared to its previous wireless generations, such as 5G [144]. A trial using 6G for V2X communications was introduced in [140], where intelligent reflecting surfaces were integrated to facilitate V2X communications by addressing signal blockage and large-scale fading due to other vehicles and ubiquitous objects in urban cities. Thus, intelligent reflecting surfaces were employed as signal relays to assist data transmissions, aiming to achieve two performance targets: latency reduction and transmission rate enhancement [145]. 6G-based intelligent reflecting surfaces were also exploited to optimize the achievable transmission rate for V2X communication links between the base station and CAVs [141]. Channel gains for vehicle-based station communications were formulated, phase-shift matrices were constructed in a fashion, and the data rate was optimized via a simplified convex optimization method. In addition to transmission rate optimization for V2X Communications, 6G offers flexible channel management solutions via Multiple-input Multiple-output (MIMO) settings, where beamforming management plays a crucial role [146]. In this regard, a channel subspace structure was constructed using received vehicle signals evaluated via perceptive ray-tracing channel data and vehicle trajectories.
Integrating 6G technology into V2X communications plays a pivotal role in ensuring reliable connectivity, a critical requirement for CP applications where timely sharing of sensory data from vehicles with servers, such as base stations, is essential. Effective resource management is paramount in guaranteeing adequate resources to all vehicles to facilitate successful and swift data transmission. This incentive led the work in [142] to propose a resource allocation strategy tailored for 6G-based V2X communications. This strategy allocates resource blocks to participating CAVs within a shared vehicular network, guided by reinforcement learning principles. This approach allows for formulating allocation decisions within an environment shaped by interactions between the base station and end users. Unlike conventional optimization methods that rely on complex mathematical formulations, reinforcement learning estimates resource allocation in an equitable and optimal manner while mitigating transmission interference, all within the constraints of radio resources.
Figure 5: Illustration of a 6G-based V2X communication system.
Figure 6: The modern generic framework of the cooperative perception for autonomous driving.
Furthermore, one of the significant challenges in CP applications revolves around potential conflicts among vehicles, particularly when they share the same wireless channels for data transmission and operate within the same coverage area. The work in [143] proposed a cooperative collision detection mechanism to address this issue. This mechanism leverages reinforcement learning techniques to establish a near-optimal policy governing vehicle trajectories and V2X communications in the context of a 6G wireless network environment. The optimization process considers factors such as latency and quality of service in accordance with 6G standards, utilizing practical datasets to evaluate collision possibilities in dense vehicle networks.
For convenience, the related works in V2X communications are summarized in Table III, highlighting each reference work’s key contribution. Fig. 5 shows an example of 6G-based V2X communication systems. Regarding the standardization of V2X communication, protocol settings and corresponding use cases, the readers can refer to [147].
IV Modern Generic Framework of Cooperative Perception
This section introduces a contemporary and comprehensive framework for CP in autonomous driving, along with relevant issues and methodologies.
IV-A Framework Overview
Given multiple agents capable of perception and communication within a shared environment, the objective is to enhance each agent’s perceptual accuracy through distributed cooperation. Fig. 6 illustrates the operational process of the proposed framework, comprising two major stages: single-agent perception and multi-agent collaboration.
During the single-agent perception phase, each agent acquires measurements of its Six Degrees-of-Freedom pose while simultaneously perceiving the surrounding environment. Communication messages are generated by packaging Cooperative Awareness Messages (CAMs) [148] and Collective Perception Messages (CPMs) [149]. CAM messages convey the agent’s status, encompassing information such as position and motion state, while CPM messages provide additional details regarding objects detected through sensor data, such as other vehicles and pedestrians. It is worth noting that the content of perceptual information may encompass raw sensor observations, intermediate features, and/or detection outcomes, contingent upon the chosen style of cooperation.
The collaboration stage encompasses three primary components, as depicted in Fig. 6: preparation for transmission, cooperative information alignment, and perception information fusion. To ensure efficient communication while preserving the accuracy of the cooperation system, the initial step involves carefully selecting suitable cooperative agents and pertinent perceptual information. Subsequently, each agent shares the CAM and the refined CPM with its peers through V2X communication. Following this, the ego vehicle aligns the acquired perception information with local data through coordinate conversion or pose transformation techniques. Finally, the aggregated data undergoes consolidation via a fusion algorithm, followed by decoding, resulting in a conclusive perception outcome. Note that all CAVs can transmit their perceptual data to a roadside infrastructure via V2X networks. The associated edge computing infrastructure is responsible for aggregating and consolidating this incoming data, which is then distributed to CAVs according to their specific needs. In the subsequent sections, we provide detailed insights and a review of existing works on the collaboration stage’s primary components.
IV-B Preparation for Transmission
TABLE IV: Summary of the Methods of Preparation for Transmission.
Category Method Year Scenario Modality Dataset/Simulator Cooperation Style Downstream Task Available Code
CAS CPOB-LearnCom [150] 2022 V2I LiDAR CARLA-3D [150] Intermediate Detection $\times$
CODS [151] 2023 V2V Camera & LiDAR CARLA Intermediate Detection $\times$
AVUCB [152] 2022 V2X Camera MS-COCO [153] Early Detection $\times$
MASS [154] 2023 V2V LiDAR DOLPHINS Early Detection $\times$
SelectComm [155] 2023 V2V LiDAR AUTOCASTSIM Intermediate Detection $\times$
IoSI-CP [156] 2023 V2X LiDAR OPV2V,V2XSet Intermediate Detection $\surd$
PIS DyDisse[157] 2019 V2X Camera/LiDAR -- Late Tracking $\times$
VehIdenti [158] 2023 V2X -- CARTERY [158] -- -- $\times$
RedunMitig [159] 2019 V2X -- SUMO Late Detection $\times$
Look-ahead [160] 2019 V2X -- SUMO Late Detection $\times$
AICP [161] 2022 V2V Camera -- Late Detection,Tracking $\surd$
DFS [162] 2023 V2X LiDAR CARTI Early&Intermediate&Late Detection $\times$
DRLCP [163] 2020 V2X Camera CIVS [163] Late -- $\times$
FS-COD [164] 2020 V2V LiDAR CARLA Intermediate Detection $\times$
COOPERNAUT [165] 2022 V2V LiDAR AUTOCASTSIM Intermediate Policy Learning $\surd$
CenterCoop [166] 2023 V2I LiDAR DAIR-V2X Intermediate Detection $\times$
Slim-FCP [167] 2022 V2V LiDAR KITTI, T&J Intermediate Detection $\times$
FPV-RCNN [168] 2022 V2V LiDAR COMAP [169] Intermediate Detection $\surd$
MRSR [170] 2023 V2X LiDAR V2X-Sim,OPV2V Intermediate Detection $\surd$
GevBEV [171] 2023 V2V LiDAR OPV2V Late Detection,Segmentation $\times$
AIS Where2comm [172] 2022 V2X Camera/LiDAR V2X-Sim,DAIR-V2X, OPV2V,CoPerception-UAVs Intermediate Detection $\surd$
  • Note: CAS: Cooperative Agent Selection, PIS: Perception Information Selection, AIS: Joint Agent & Information Selection.
In CP systems relying on V2X communication, the effective deployment of diverse sensors and continuous dissemination of information across multiple agents are essential to provide comprehensive traffic awareness. However, disseminating data without appropriate information filtering can result in substantial communication and computational resource usage, along with a high volume of less pertinent data. For example, neighbouring vehicles might redundantly transmit information about the same objects, causing inefficient utilization of communication channels [173]. This excessive network load elevates the risk of important data packets experiencing delays or being lost, potentially compromising CP performance and safety. To actualize the potential of V2X CP, it is essential to prioritize a swift, secure, and reliable flow of information between vehicles and associated entities like road infrastructures, users, and network nodes, especially given the unpredictable and rapidly changing nature of traffic situations [174].
To avoid communication congestion issues, several researchers have focused on reducing shared data volume before transmission, particularly when bandwidth among agents is limited. As summarized in Table IV, existing strategies aim to alleviate the load on the vehicular network while maintaining the performance of the CP system. These strategies fall into three categories:
  • Cooperative agent selection: This category focuses on minimizing communication costs by transmitting perception information over a partially connected agent graph, thus reducing redundancy;
  • Perception information selection: This category addresses the maximization of informativeness, where cooperative agents select a subset of perceptive data crucial to the ego-vehicle, optimizing the data shared;
  • Agent & information selection: This category encompasses a joint consideration of both cooperative agent selection and perception information selection, aiming to balance communication efficiency and data importance.
IV-B1 Cooperative Agent Selection
Selecting the right collaborators is crucial because inappropriate candidates can cause disruptions and disturbances that could harm overall performance. In recent years, communication strategies have been created to help choose beneficial partners for cooperation, considering perception improvement and communication quality.
The attention mechanism, recognized for its efficacy in improving model efficiency by focusing on relevant data, offers promise in selecting suitable agents for the ego CAV. In this context, an attention-based communication scheme was introduced in [150] for V2I cooperative object detection. This scheme leverages a general attention mechanism, where each infrastructure calculates a matching score based on local and received information within their cooperative system. Subsequently, they broadcast this score along with CPMs to the collaborative vehicle. Upon receiving data packets from connected infrastructures, the ego vehicle employs an attention mechanism to select a neighbouring infrastructure for cooperation. While this approach conserves transmission bandwidth by choosing only one agent for V2I communication, it does not harness global information optimally.
Multi-agent Cooperation can be modelled as a topology graph where each node represents a connected agent and transmission messages are generated and broadcasted periodically. In [151], a graph-based on-demand data sharing mechanism was proposed to address network constraints. The feature maps from each vehicle are compressed through quantization to reduce communication costs. Ego-vehicles receive feature maps only from selected coop-vehicles located within their perception range, establishing communication channels and continuously receiving the message body. The on-demand sharing mechanism enabled efficient cooperation among connected vehicles by selectively sharing compressed feature maps but ignored the integration of spatial-temporal information.
To minimize energy consumption, the sensor sharing scheduling problem can be framed as a variant of the Multi-Armed Bandit (MAB) problem [175], factoring in perception performance, time-varying wireless channels, and power consumption. In this regard, an online learning-based algorithm known as the Adaptive Volatile Upper Confidence Bound (AVUCB) algorithm was proposed by reference [152]. This algorithm schedules the most advantageous vehicle, offering an enhanced view while minimizing total power consumption and adhering to communication bandwidth and detection accuracy constraints. However, the AVUCB algorithm does not account for the high mobility of cooperative vehicles. Addressing this mobility challenge, the Mobility-Aware Sensor Scheduling (MASS) algorithm [154], an extension of the AVUCB, formulated decentralized CP scheduling as a Restless-MAB (RMAB) problem [176]. In the RMAB context, rewards continually evolve, requiring each vehicle to continuously learn about its surroundings while selecting actions that empirically maximize rewards. When applied to a practical SUMO trace, the MASS algorithm demonstrated the ability to enhance overall perception gain without incurring additional costs associated with frequent meta-information exchanges.
Cooperative agent selection can also be likened to a Multi-agent Path Finding (MAPF) problem, which plans paths for multiple agents [177]. Inspired by MAPF, a selective communication algorithm [155] was proposed to determine the subset of cooperative vehicles based on information gains from connected vehicles. Each vehicle predicts the set of objects in its field of view and transmits the centre locations of all detected objects to the ego-vehicle. Upon receiving this location information from all vehicles, the ego-vehicle estimates information gains by comparing its detected objects’ positions to received 2D centre positions from other connected cars. The vehicles associated with the highest information gain are selected to form the communication scope. In contrast to random selection, this algorithm achieved higher success rates with minimal additional communication costs.
However, all the aforementioned methods primarily focused on spatial correlations, overlooking the significance of temporal dependencies in collaborator selection. To address this gap, IoSI-CP [156] capitalized on the importance of semantic information from both temporal and spatial dimensions to select optimal collaborators. This approach leveraged a Graph Neural Network (GNN) to capture inherent correlations between the ego vehicle and its neighbouring agents. Each agent was represented as a node, encompassing sparse feature maps and latency information. The GNN assigned weight values to each agent with respect to the ego vehicle. Agents with weights of zero or lower were considered ineligible for collaboration. This refined weighting effectively eliminated unsuitable collaborators responsible for noise, leading to improved perception performance. Nevertheless, this approach relied on one-dimensional weights, which may not fully capture the intricacies addressed by high-dimensional representation methods.
IV-B2 Perception Information Selection
Recently, there has been a growing concern regarding the selection of cooperative message content. A number of methods have emerged for selecting or generating CPMs intended for transmission. These methods can be broadly categorized into rule-based, distance-based, and learning-based approaches, each rooted in distinct mechanisms and theoretical principles.
The rule-based approach to CPM generation/selection entails designing rules that dictate which aspects of detected object information should be included in a CPM. For example, the work in [157] proposed a rule that transmits information about an object’s mode (including acceleration and yaw rate) only if the object is turning or accelerating. Similarly, in [158], visual information from detected vehicles, such as license plates, vehicle brands, and colours, was used to determine their communication IDs. These IDs are then used to integrate CPMs with local perception information, ensuring the confidentiality of the V2X architecture. All detected vehicles that have exchanged CAM messages will not be considered in future CPMs. Standardization efforts, led by the European Telecommunications Standards Institute (ETSI), have also introduced CPM generation rules, specifying what information CPMs should include and when to generate new messages. Further, the work in [159] and [160] has proposed improved algorithms for these rules to filter out redundant information about detected objects that nearby vehicles have recently transmitted. While rule-based methods are straightforward to implement, they rely primarily on changes in the position and speed of detected objects.
Distance-based methods prioritize data based on distance metrics, considering the spatial location relative to the ego vehicle’s onboard sensors to measure data point significance. For instance, the Augmented Informative Collaborative Perception (AICP) system [161] employed a fitness sorting algorithm based on Mahalanobis distance to select crucial information at the application level, effectively filtering out less relevant packets with minimal delay. In a separate approach, Dynamic Feature Sharing (DFS) [162] addresses limited communication bandwidth challenges by identifying relevant feature data for sharing through Random-K priorities based on Manhattan distance. DFS incorporates early, intermediate, and late fusion strategies within the CP system to optimize dynamic node engagement.
Rule-based methods primarily select specific attributes of detected objects for dissemination, while distance-based methods emphasize the transmission of feature cells or data points close to the ego vehicle. However, these approaches may only partially meet the ego vehicles’ perception information requirements in real-world scenarios, leading to suboptimal results. In contrast, learning-based approaches leverage the entirety of perception data to intelligently determine information selection for sharing.
Deep Reinforcement Learning (DRL) allows agents to learn policies through interactions with their environment, maximizing cumulative rewards [178, 179, 180]. DRL techniques can be applied to vehicular communications to optimize data selection for transmission. The work [163] introduced a DRL-based Cooperative Perception (DRLCP) scheme, where each CAV intelligently decides what information to transmit based on the context captured by its onboard sensors. DRLCP’s primary objective is to determine whether to transmit or discard data, allocating communication resources to critical packets and reducing redundancy. However, it was evaluated on only two road networks without considering the impact of obstructions on vehicular communications.
Sparse Bird’s Eye View (BEV) features extracted from point clouds can be further compressed to minimize redundancy by extracting valuable semantic information. Feature encoding with Convolutional Neural Networks (CNNs) is a widely adopted learning-based method for reducing CPM size. For instance, FS-COD [164] modified the backbone structure of Complex-YOLO [181] based on BEV image characteristics. Filters at the last layer of the CNN based feature extractor determine the size of shared data between cooperative vehicles. COOPERNAUT [165] designed a Point Encoder using the Point Transformer [182] for each connected vehicle to learn LiDAR information representations. This approach processes and encodes point clouds into key points with compact representations learned by Point Transformer blocks, filtering partially processed data to balance perception accuracy and communication load. CenterCoop [166] encoded the informative cues from the whole BEV context into compact center representations. The communication cost can be significantly reduced by transmitting compact center features as queries instead of the whole feature map.
While CNN-generated feature maps offer potential data volume reduction, their dimensions often exceed transmission feasibility within current V2X technology scopes [167]. To address this concern, a comprehensive strategy was adopted in [167], combining feature encoding, channel selection, and data compression techniques to optimize feature maps. Specifically, an autoencoder architecture, including a channel-wise semantic information encoder and a pointwise convolution decoder, optimized the representation of sparse feature maps. The SENet channel attention module [183] weighed channel importance, determining optimal channels for transmission based on attention weights and the uniqueness of semantic information. A learning-based Irrelevant Feature Remover (IFR) further reduced the transmitted feature map size by replacing extraneous pixels with zeros. This approach significantly reduced data transmission size compared to solutions relying solely on feature encoding.
Another category of learning-based approaches, the selective-region mechanism, focuses on selecting CPM information by considering relevant spatial areas. These methods ensure that only pertinent information within appropriate communication regions is transmitted, avoiding bandwidth waste. For instance, [168] employed a bounding box matching module and strategic critical point selection to choose features enclosed within bounding box proposals for transmission. In contrast, MRSR [170] introduced the Entropy-CS, a trainable region-wise communication selection module, leveraging entropy to distinguish informative regions and select suitable regions for transmission based on resolution levels. GevBEV [171] identifies areas requiring supplemental information for the ego by utilizing evidential BEV maps. This method employs evidential deep learning, incorporating Dirichlet evidence [184], to quantify point-based classification uncertainty within BEV maps. This quantification aids in pinpointing the most relevant regions for communication.
TABLE V: Summary of the Works of Temporal and Spatial Alignment.
Category Method Year Scenario Modality Dataset/Simulator Cooperation Style Downstream Task Available Code
ACD FFNet [185] 2023 V2I LiDAR DAIR-V2X Intermediate Detection $\surd$
SyncNet [186] 2022 V2V LiDAR V2V-Sim Intermediate Detection $\surd$
V2VNet [187] 2020 V2V LiDAR V2V-Sim Intermediate Detection, Motion Prediction $\surd$
CoBEVFlow [188] 2023 V2V LiDAR CARLA, DAIR-V2X Intermediate Detection $\surd$
APE RANSAC-based [189] 2022 V2V LiDAR COMAP Late Detection, Location $\times$
RobustV2VNet [190] 2021 V2V LiDAR V2V-Sim Intermediate Detection, Motion Prediction $\times$
CoAlign [191] 2022 V2X LiDAR OPV2V,V2X-Sim,DAIR-V2X Intermediate Detection $\surd$
ICP&OP [192] 2022 V2V LiDAR OPV2V Late Detection $\times$
Map Container [193] 2022 V2X Camera,LiDAR VISSIM Late State Estimation $\times$
SCOPE [194] 2023 V2X LiDAR OPV2V,V2XSet,DAIR-V2X Intermediate Detection $\surd$
CBM [195] 2023 V2V Camera,LiDAR SIND[196], OPV2V Late Detection $\surd$
DMGM [197] 2023 V2V Camera CARLA & SUMO Early -- $\surd$
ACD & APE V2X-ViT [198] 2022 V2X LiDAR V2XSet Intermediate Detection $\surd$
  • Note: ACD: Alignment for Communication Delay, APE: Alignment for Pose Errors.
IV-B3 Joint Agent and Information Selection
Methods that consider both cooperative agent and information selection can offer a more balanced approach to optimizing precision and bandwidth in CP. For example, Where2comm [172] introduced a spatial-confidence-aware CP framework that utilizes spatial confidence maps generated for each agent to identify perceptually critical areas. These critical features are then compactly packaged as messages and shared with others via a sparsely connected communication graph. A communication graph construction strategy is employed to determine with whom to communicate to minimize unnecessary bandwidth usage. After communication, confidence-aware multi-head attention is applied to enhance each agent’s feature map by incorporating received messages from other agents. Experimental results demonstrate that Where2comm efficiently enables mutually beneficial cooperation by sharing spatially sparse yet critical features in V2X communication.
IV-C Cooperative Information Alignment
To enable precise information fusion in CP, it is crucial to temporally and spatially align the perceptual information received via V2X communication with the ego data [199]. While many existing methods aim to synchronize exchanged messages within a consistent spatial coordinate system to address this challenge, practical issues such as misalignment and inconsistency can arise due to communication delays and pose noise [200]. Therefore, mitigating the effects of latency and pose errors is essential rather than relying solely on direct synchronization of asynchronous information. In this subsection, we provide a detailed discussion of all the aspects outlined in Table V.
IV-C1 Alignment for Communication Delays
In practical V2X communication scenarios, latency issues are inevitable due to various real-world factors, including limitations in wireless communication, model complexity, and the number of agents involved [201]. The presence of communication latency can mislead detectors, resulting in significant deviations in perception results and ultimately causing fusion misalignment. Temporal asynchrony can detrimentally affect the performance of the CP system, potentially leading to outcomes worse than those of single-vehicle perception [202]. Hence, developing a robust CP system capable of handling unavoidable communication delays is imperative.
A common and effective strategy to address the latency issue is to predict the sender’s perception at the current time step and subsequently perform coordinate transformations using the predicted data frame. In this context, the FFNet approach [185] exploits the temporal coherence of sequential infrastructure frames to predict future features and counteract asynchrony. FFNet adopts a similar architectural framework to that of SECOND [203], which includes a Backbone and Feature Pyramid Network, to extract the first-order derivative of feature flow from historical infrastructure frames. However, it is important to note that this first-order expansion method is most effective when the time interval is very short.
An alternative approach, proposed in [186], introduces a latency compensation module called SyncNet. SyncNet utilizes time series prediction techniques to estimate received features at the current time step based on historical information. To achieve this, a dual-branch pyramid Long Short-term Memory (LSTM) network is employed to infer real-time features and corresponding cooperative attention. Additionally, a time modulation module is integrated to allocate temporal attention between the estimated cooperative feature and the received asynchronous features, taking into account the latency time. The final estimation is generated by combining the estimated feature with the latest asynchronous feature based on the temporal weight.
Another approach for mitigating the impact of communication delays involves re-encoding received features with delay time. For instance, V2X-ViT [198] addresses time delay through a Delay-aware Positional Encoding (DPE) module that utilizes sinusoid functions conditioned on delay time and channel information to initialize a learnable embedding. This embedding is further updated through a linear projection. The resulting projected embedding, serving as an extension of temporal delay, is integrated into the received feature, thus achieving temporal alignment of features. In the case of V2VNet [187], each vehicle employs a fully connected GNN as an aggregation module, where each node represents the state of a CAV. Within this GNN, a CNN is utilized to compensate for latency, taking inputs such as the received intermediate feature, relative 6DoF pose, and delay time. The outcome is a time-delay-compensated representation concatenated with zeros to enhance the node’s ability to aggregate information from other vehicles.
Previous works have not addressed the issues raised by common irregular time delays, rendering existing collaborative perception systems unable to reach their full potential in real-world scenarios. As a substitute, CoBEVFlow [188] considered each collaboration message as one irregular sample and used irregular time series forecasting to estimate the BEV flow map for asynchronous feature alignment. In this way, the historical frames were used to estimate a BEV flow map, which encodes the motion information in each grid cell. The asynchronous perceptual features can be reassigned to appropriate spatial locations based on the BEV flow map and region of interest generation, thereby aligning perceptual features on the time dimension and mitigating the impact caused by asynchrony.
IV-C2 Alignment for Pose Errors
Establishing precise estimates of relative transformation matrices is foundational for V2X cooperation, ensuring consistent alignment of received perception information in a shared spatial coordinate system. In practical scenarios, these transformation matrices rely on sensor measurements from posture detection systems, such as the Global Positioning System (GPS), Real-Time Kinematic (RTK), and Inertial Measurement Unit (IMU). However, real-world systems are prone to imperfections, including noise and limited accuracy, which can introduce errors into the transformation process, potentially degrading perception performance.
Efforts have been made to address these challenges using statistical learning models. For example, one approach [189] corrected relative locations between two CAVs by matching key points in their perception results with the Random Sample Consensus (RANSAC) algorithm [204]. This method used the consensus of anchor points, such as vehicle centres and static object points, to calculate relative pose transformations accurately. While effective, this approach required the setting of problem-specific thresholds.
Within the framework of V2VNet [187], a consistency module [190] was introduced to improve relative pose estimations. This module structured the global coherence of absolute poses as a Markov random field, incorporating Bayesian reweighting. However, it is important to note that this model relies on ground-truth poses during the training phase, which may be limited in practical scenarios. In contrast, CoAlign [191] proposed an Agent-Object Pose Graph Optimization method for enhancing pose consistency without requiring precise supervision during training. This approach utilized an agent-object pose graph to model relationships and applied graph optimization techniques to achieve accurate pose alignment. Notably, CoAlign assumed that pose noise followed a Gaussian distribution, a simplifying assumption to facilitate the process.
The Iterative Closest Point (ICP) [205] algorithm is widely used for aligning 3D point clouds. In [192], ICP was employed to align two sets of points, where one set represented 3D bounding boxes detected by the Ego, and the other set represented those detected by the sender. This source and target point alignment were achieved using an ICP-based matching algorithm, iteratively optimizing the pose transformation to achieve a locally optimal alignment. However, it is important to note that ICP can be sensitive to points lacking correspondences, and its neighbour selection step relies on a parameter that directly impacts the registration outcome. In response to this challenge, the relative transformation was addressed as an optimal transport problem [206], seeking to minimize transportation costs between the source and target points. Nevertheless, the effectiveness of these approaches is contingent on the number of targets present within the shared FoV of both the Ego and the sender.
High-definition (HD) maps represent a valuable resource for achieving precise spatial alignment, primarily due to their ability to provide accurate self-vehicle localization. In this regard, a map-based CP framework called the “map container” was introduced in [193]. This framework utilizes the HD map to establish precise benchmarks for spatiotemporal transformations of perceptual data. The core concept involves automatically transforming multi-agent perceptual information into the map coordinate space through a map-matching algorithm. This integration process uses diverse information sources within a decentralized fusion framework. The map container directly incorporates received information into the joint state estimation optimization process, showcasing a robust fusion architecture.
Utilizing multi-scale feature interaction can effectively mitigate feature map misalignment arising from pose errors. This approach involves encoding the feature map into multiple scales and subsequently aggregating them to derive the ultimate collaboration feature. To address this concern, a spatiotemporal awareness CP framework known as SCOPE was introduced in [194]. SCOPE aims to obtain robust representations while accounting for localization errors, achieved through confidence-aware multi-scale feature interaction. Specifically, the framework employs the spatial-confidence-aware method [172] to select pertinent features and generate confidence maps. These confidence maps are further encoded into three distinct scales. A pivotal component of SCOPE is the application of a deformable cross-attention layer [207], which facilitates the aggregation of information from key points extracted from the confidence maps. This aggregation process is governed by a threshold-based selection function applied at each scale. The final collaboration feature is generated by amalgamating information derived from these three scales through the utilization of a convolutional layer.
In the context of addressing the impact of GPS localization inaccuracies, the Multi-scale Window Attention Module (MSWin) introduced in V2X-ViT [198] plays a crucial role. MSWin is designed to learn long-range spatial interactions at various scales through a pyramid of resolution windows. Each of these windows accommodates a different attention range, enabling the simultaneous capture of both local and global spatial feature interactions. This approach enhances detection robustness in the face of positioning errors.
TABLE VI: Summary of the Works of Cooperative Information Fusion.
Method Year Scenario Modality Dataset/Simulator Cooperation Style Downstream Task Available Code
Cooper [111] 2019 V2V LiDAR KITTI, T&J Early Detection $\times$
F-Cooper [208] 2019 V2V LiDAR KITTI, T&J Intermediate Detection $\times$
PillarGrid [209] 2022 V2X LiDAR CARTI Intermediate Detection $\times$
CoCa3D [210] 2023 V2X Camera OPV2V$+$, DAIR-V2X, CoPerception-UAVs$+$ Intermediate Detection $\surd$
CPM-LCF [211] 2021 V2X LiDAR -- Late Detection $\times$
CI-based [212] 2022 V2X LiDAR CARLA Late Tracking $\times$
V2VNet [187] 2020 V2V LiDAR V2V-Sim Intermediate Detection, Motion Prediction $\surd$
DiscoNet [213] 2021 V2V LiDAR V2X-Sim Intermediate Detection $\surd$
V2X-ViT [198] 2022 V2X LiDAR V2XSet Intermediate Detection $\surd$
HYDRO-3D [214] 2023 V2X LiDAR V2XSet Intermediate Detection $\times$
CoBEVT [215] 2022 V2V Camera OPV2V Intermediate Segmentation $\surd$
COOPERANAUT [165] 2022 V2V LiDAR AUTOCASTSIM Intermediate Policy Learning $\surd$
IoSI-CP [156] 2023 V2X LiDAR OPV2V, V2XSet Intermediate Detection $\surd$
Coop3D-Infra [110] 2022 V2I LiDAR CARLA Hybrid Detection $\surd$
ML-Cooper [216] 2022 V2V LiDAR KITTI Hybrid Detection $\times$
Recently, an innovative approach called inter-agent object association has emerged as a solution to mitigate localization errors in pose transformation. This method involves associating key points or objects within observations provided by connected agents, determining optimal relative pose transformation based on matched correspondences. In one approach presented in [195], a Context-based Matching (CBM) algorithm was introduced for inter-agent object association. CBM involves generating object clusters, considering each bounding box’s heading direction in local coordinate systems. Object correspondences are established through an iterative pairwise comparison procedure applied to all clusters to maximize global consensus. However, identifying correspondences can be challenging due to visual folding, non-visible objects, and noisy perception. To overcome these challenges, another method proposed in [197] introduces a deep masked graph-matching approach that frames correspondence identification as a graph-matching problem. This approach employs a masked GNN and includes a mechanism for handling non-visible objects. The network integrates multi-modal visual, spatial, and GPS information using SoftMax variance thresholding. Each node in the graph represents an object, while edges capture the spatial relationships between pairs of objects. This deep masked neural network explicitly addresses challenges related to object non-visibility caused by occlusion and limited field of view, ultimately enhancing the precision of correspondence identification.
IV-D Cooperative Information Fusion
The effectiveness of CP crucially depends on the design of an efficient fusion strategy, which aims to consolidate the shared information to provide a coherent and accurate understanding of traffic situations. The following will examine recent fusion methods employed in CP information. For reference, Table VI provides a summary of all discussed fusion models in this subsection.
IV-D1 Traditional Fusion
Primary reduction operators like summation, maximum, or average pooling have found extensive use in cooperative information fusion for multi-agent perception. For instance, in [111], a raw-data CP system, dubbed Cooper system, was introduced for CAVs to enhance 3D object detection. However, this early cooperative approach entailed a substantial data overhead for transmitting raw point cloud data. F-Cooper [208] extended Cooper to the feature level to balance bandwidth constraints and detection accuracy in cooperative 3D object detection. Deep features were acquired through neural networks at each vehicle’s side, and max pooling was employed to combine V2V voxel features. Another model, PillarGrid [209], operated at the feature level for cooperative 3D object detection based on point cloud data. It employed a grid-wise feature fusion approach using max-pooling to integrate deep features from multiple agents. CoCa3D [210] leveraged point-wise maximum fusion to amalgamate BEV features from several CAVs to enhance camera-based detection through cooperative depth estimation. These CP methods outperformed single-vehicle detection techniques regarding 3D object detection performance.
In [211], a rule-based fusion scheme for CPMs was proposed. It identified conditional branches for associating locally detected objects with those detected via V2X communication. This fusion scheme was designed for scenarios where CAVs generated CPMs following standardized ETSI rules and transmitted them to a fusion centre via wireless communication. The fusion centre applied the proposed algorithm to generate merged CPMs, which were subsequently broadcast to all CAVs in range for downstream perception tasks. Simulation results at an intelligent junction illustrated the relevance and effectiveness of this fusion strategy for obstacle detection.
Non-parametric max operators and rule-based CPM fusion can integrate information with minimal computational cost but may fail to identify data correlations. Parametric methods, on the other hand, often yield optimal estimates by mathematically incorporating measurements and their associated uncertainties, resulting in more accurate and reliable fusion outcomes. For instance, the Covariance Intersection (CI) algorithm can consistently fuse estimates without prior knowledge of cross-correlations between them. Building upon this concept, a CI-based fusion strategy for CP was introduced [212]. This approach achieves a consistent fusion of local perception sensor data and perceived object information from other agents through V2X communication. However, it is important to note that high-level fusion methods, like the one described above, can be bandwidth-efficient. Still, they may encounter challenges when dealing with potentially noisy and incomplete single-agent perceptions, leading to less than satisfactory fusion results.
IV-D2 Graph-based Fusion
Multi-agent collaboration can be effectively represented as a collaboration graph, where nodes represent individual agent states, and edges signify pairwise collaboration between agents. At the core of this collaboration graph lies the GNN, a vital component capable of extracting relational data representations and generating valuable node embeddings. These embeddings are derived from local node attributes and encompass information from neighbouring nodes within the graph. Consequently, node states can be continuously updated through inter-node communication and a neural network aggregation mechanism that combines received information.
In V2VNet [187], each vehicle employed a fully connected GNN [217] as an aggregation module, enabling intelligent data fusion from various viewpoints. Specifically, received messages were aggregated at each node using a mask-aware accumulation operator, ensuring consideration only of overlapping FoVs. Subsequently, a Convolutional Gated Recurrent Unit (ConvGRU) was utilized to update the state of each node and generate the fused intermediate representation. The gating mechanism of ConvGRU facilitated information selection for the accumulated received messages based on the current belief of the receiving CAV. The study showed that V2VNet successfully balanced accuracy improvements and bandwidth requirements. However, it employed a scalar-valued edge weight to indicate the importance of the neural message from one agent to another, which could not effectively model the strength of collaboration between two agents.
On the other hand, DiscoNet [213] employed an edge encoder to establish a correlation between the ego feature map and the feature map from another agent, generating a matrix-valued edge weight. The edge encoder concatenated two feature maps along the channel dimension and applied four convolutional layers for downsampling. Additionally, a softmax operation was performed at each cell in the feature map to normalize the edge weights across multiple agents. Furthermore, knowledge distillation was utilized to enhance training by constraining post-cooperation feature maps in the intermediate fusion-based student model to align with the early fusion-based teacher model’s correspondences. This approach promoted improved feature abstraction and aggregation, enhancing object detection performance.
IV-D3 Attention-based Fusion
The attention mechanism introduced in the transformer model has found widespread use in determining fusion weights. Building on the concept of the heterogeneous graph transformer [218], the study presented in [198] introduced a V2X-based Vision Transformer (V2X-ViT). This model incorporates two key modules: the Heterogeneous Multi-agent Self-attention (HMSA) module and the Multi-scale Window (MSwin) attention module. The HMSA module discerns unique characteristics within the sensor measurements obtained from infrastructure and CAVs. In the HMSA module, agent types are associated with nodes and edges within a directed graph. Importance weights between pairs of nodes are computed based on the corresponding node and edge types. The MSwin module employs multiple window sizes to aggregate spatial information, thereby enhancing detection robustness in the presence of localization errors. Several V2X-ViT blocks are stacked to iteratively learn inter-agent interactions and per-agent spatial attention, ultimately yielding an aggregated feature representation for detection tasks. To further enhance object detection performance, HYDRO-3D [214] extends the V2X-ViT framework to create a hybrid object detection and tracking CP model. This extension integrates V2X-ViT’s object detection features with historical object tracking information. Object tracking features are extracted from previously tracked objects using a spatial-temporal pyramidal 3D network. Subsequently, an operation concatenates the object detection and tracking features along their channel dimension to enable robust object inferences. Experimental results demonstrate that HYDRO-3D leverages longer-term historical information from object tracking to enhance object detection accuracy.
LiDAR sensors offer precise 3D depth measurements but are costly to implement at scale. On the other hand, camera-based object detection presents a more cost-effective solution with a straightforward setup for object localization. Recognizing this advantage, the work in [215] introduced a sparse vision transformer framework named CoBEVT. This framework leverages features derived from multi-agent multi-camera sensors to collaboratively generate BEV-segmented maps. Within CoBEVT, a 3D vision transformer named FuseBEVT was introduced to attentively fuse information obtained from BEV features shared among multiple agents. A pivotal component of FuseBEVT is the Fused Axial (FAX) attention module, which facilitates sparse local and global spatial interactions across various multi-agent and multi-camera perspectives. FAX integrates a global attention mechanism to capture long-range relationships and a local attention mechanism aggregating intricate regional characteristics. CoBEVT exhibited superior performance in BEV segmentation tasks by exploiting the FAX module’s effectiveness in feature fusion.
The Point Transformer [182] is a highly effective method for learning compact point-based representations from 3D point clouds. It excels at capturing non-local interactions among points and generating permutation-invariant representations, rendering it particularly adept at aggregating point clouds from multiple vehicles. In the context of COOPERNAUT [165], a point transformer-based Representation Aggregator (RA) was introduced for information fusion purposes. This RA consists of two main components: a voxel max-pooling operation and a point transformer block. The voxel max-pooling operation is employed to spatially aggregate incoming messages close by down-sampling all the points within the same voxel grid cell. Subsequently, the point transformer block is used to fuse the multi-agent perception information. Empirical results have demonstrated that COOPERNAUT significantly enhances the robustness of self-driving strategies, particularly in scenarios involving risk-sensitive traffic situations.
The attention-based models mentioned earlier have made significant strides in cooperative information fusion, primarily focusing on spatial dimensions. However, they often operate on single frames, which can lead to difficulty detecting fast-moving objects, given the sparsity of point clouds. Additionally, the shared information may suffer from asynchrony and misalignment in both spatial and temporal dimensions due to transmission delays and sensor configuration disparities. Thus, it becomes imperative to concurrently address spatial and temporal aspects of shared information to achieve accurate perception.
To address the above challenges, SCOPE [194] introduced a Context-aware Information Aggregation (CIA) approach to incorporate spatiotemporal information from previous frames of the ego agent. A pivotal component of CIA is a pyramid LSTM, which replaces the conventional matrix multiplication of LSTM with a multi-scale convolutional structure. The resultant hidden state from the pyramid LSTM serves as a refined context-aware feature. This feature and collaboration and ego features are fused based on their complementary contributions and spatial attention maps obtained through max pooling. However, it is worth noting that max-pooling may not fully capture the underlying correlations among informative features. In contrast, IoSP-CP [156] devised a Historical Prior Hybrid Attention (HPHA) fusion algorithm that leverages a multi-scale transformer and a short-term attention module. HPHA assigns varying weights to different information from collaborators to optimize feature fusion. The multi-scale transformer module extracts attention weights at different scales, facilitating collaborative correlation learning across distinct spatial dimensions. The short-term attention mechanism captures temporal dependencies and identifies discriminative object features from the temporal dimension. HPHA empowers IoSP-CP to enhance informative features while reducing background noise.
TABLE VII: Summary of the Works on Issues of Cooperative Perception and Corresponding Methods.
Category Method Year Scenario Modality Dataset/Simulator Cooperation style Downstream task Available Code
MH MAMP [219] 2022 V2V LiDAR OPV2V Late Detection $\surd$
UDT2T [220] 2022 V2V -- CARLA Late Tracking $\times$
MPDA [221] 2022 V2X LiDAR V2XSet Intermediate Detection $\surd$
MACP [222] 2023 V2V LiDAR V2V4Real, OPV2V Intermediate Detection $\surd$
DH HM-ViT [223] 2023 V2V Camera&LiDAR OPV2V Intermediate Detection $\surd$
DI-V2X [224] 2023 V2X LiDAR DAIR-V2X, V2XSet Intermediate Detection $\surd$
DG-CoPerception [225] 2023 V2V Camera OPV2V Intermediate Segmentation $\surd$
VINet [226] 2022 V2X LiDAR CARTI Intermediate Detection $\times$
DRR MWV2IComm[227] 2021 V2I -- -- -- Passing intersection $\times$
LC V2VAM$+$LCRN [228] 2022 V2V LiDAR OPV2V Intermediate Detection $\times$
V2X-INCOP [229] 2023 V2X LiDAR V2X-Sim,OPV2V,DAIR-V2X Intermediate Detection $\times$
DT TrustEsti [230] 2019 V2V Camera/LiDAR -- -- -- $\times$
FDII [231] 2023 V2V LiDAR CARLA Early Detection $\times$
CAD [232] 2023 V2V LiDAR Adv-OPV2V, Adv-MCity Early/Intermediate Detection $\surd$
MADE [233] 2023 V2V LiDAR V2X-sim, DAIR-V2X Intermediate Detection $\times$
PU CoPEM [234] 2022 V2X -- ViSTA Late Detection $\times$
DMMCP [235] 2023 V2X Camera/LiDAR -- Late Detection, Tracking $\times$
GevBEV [171] 2023 V2V LiDAR OPV2V Late Detection, Segmentation $\times$
Double-M [236] 2023 V2X LiDAR V2X-Sim Early/Intermediate/Late Detection $\surd$
MOT-CUP [237] 2023 V2X LiDAR V2X-Sim Late Detection, Tracking $\times$
DMSTrack [238] 2024 V2V LiDAR V2V4Real Late Tracking $\times$
TD STAR [239] 2022 V2X LiDAR V2X-Sim Intermediate Scene completion, Detection,Segmentation $\surd$
CORE [240] 2023 V2V LiDAR OPV2V Intermediate Scene completion, Detection,Segmentation $\surd$
DP FedAD [241] 2023 V2I Camera Udacity -- Turning Angle Prediction $\times$
FedBEVT [242] 2023 V2X Camera CARLA&OpenCDA -- Detection $\surd$
FL-CCS [243] 2023 V2X -- -- -- -- $\times$
S2R S2R-ViT [244] 2023 V2V LiDAR OPV2V,V2V4Real Intermediate Detection $\times$
DUSA [245] 2023 V2X LiDAR V2XSet, DAIR-V2X Intermediate Detection $\times$
  • Note: MH: Model Heterogeneity, DH: Data Heterogeneity, DRR: Data Rate Requirements, LC: Lossy Communication, DT: Data Trustworthiness, PU: Perception Uncertainty, TD: Task Discrepancy, DP: Data Privacy, S2R: Simulation to Reality.
IV-D4 Hybrid Fusion
Different collaboration schemes have advantages and limitations and can complement each other effectively. However, single-type fusion methods may not fully exploit the potential for deep fusion of various data types and generating higher-level collaborative representations. Hybrid fusion-based approaches offer promise in harnessing the strengths of diverse collaboration schemes by integrating multiple fusion methods within a single CP model.
Typically, objects near the onboard sensor exhibit a high point density and are more likely to be detected using data from a single sensor. Conversely, distant points are often sparse due to limited sensor visibility. Therefore, employing an early fusion scheme in areas far from the Ego can enhance object detection likelihood. This concept has been validated by research [110], which introduced a V2I cooperative 3D object detection model using a hybrid fusion approach to aggregate point cloud data from multiple infrastructure sensors. The central idea behind this hybrid fusion strategy is to transmit raw data (early fusion) in regions distant from a sensor where point cloud density is low and to share object-level information (late fusion) in the immediate vicinity of a sensor. This approach proves cost-effective for autonomous driving systems in complex road segments but may lack flexibility in adjusting the proportion of raw and object-level data.
In contrast, ML-Cooper [216] employed an improved Soft Actor-critic (K-SAC) [246] algorithm to dynamically reduce channel congestion. It combines three distinct cooperation schemes to maximize V2V network bandwidth utilization. Specifically, each CAV partitions its sensing data into three segments based on point cloud angle and density and shares raw, feature, and object data from each segment with other CAVs. The K-SAC algorithm determines the distribution among the three segments based on the prevailing channel conditions. This adaptability characteristic equips ML-Cooper to operate effectively in highly dynamic vehicular networks.
V Issues of Cooperative Perception and Corresponding Methods
While the majority of research on CP for driving automation has focused on key aspects of the collaboration process, such as communication latency, information alignment, and fusion strategies, as discussed in the previous section, limited attention has been given to subtle yet crucial challenges that have practical implications for advancing Cooperative ITS. This section delves into exploring recent findings and the corresponding methodologies to address these practical challenges. Table VII provides a comprehensive overview of characteristics related to these advancements.
V-A Agent Heterogeneity
Although CP systems have made significant progress, it is important to acknowledge that these improvements often assume that the agents involved are uniform. In reality, it is difficult to achieve uniformity in agent configuration and sensor equipment across connected agents, including CAVs and infrastructures, due to varying preferences among system developers and automobile manufacturers. Therefore, it is necessary to prioritize finding solutions to the practical challenges of enabling effective collaboration among diverse agents within V2X systems.
V-A1 Model Heterogeneity
Sharing perception models with identical parameters among agents could raise concerns about privacy and confidentiality, especially when CAVs come from various automotive manufacturers. Even within a single automotive enterprise, variations in vehicle types and model update frequency can result in differences in perception among model versions. However, using different models for individual agents can create domain gaps in shared perception data. Neglecting model heterogeneity can significantly reduce the effectiveness of CP, thereby diminishing the benefits of multi-agent cooperation.
Object-level-based cooperative detection frameworks can circumvent the necessity for comprehensive model data exchange by exchanging perception outcomes, including bounding boxes and confidence scores, among participating agents [219]. Nevertheless, due to model heterogeneity, the confidence scores provided by different agents may not always align accurately. In [219], a model-agnostic CP framework was proposed to tackle this confidence misalignment problem. This framework employed a customized confidence calibrator called Doubly Bounded Scaling (DBS) and a tailored Promote-suppress Aggregation (PSA) algorithm. DBS learned a scaling function based on the Kumaraswamy Cumulative Density Function [247] to eliminate confidence score bias. At the same time, PSA selected high-score bounding boxes based on calibrated confidence and spatial consistency of the same object for bounding box fusion. This approach underscored the importance of output calibration across agents with different perception models for effective late collaboration.
In addition to variations in confidence levels, connected agents’ detection outcomes may contain different information due to differences in their motion prediction models. For instance, some agents may not provide information about detected objects’ acceleration and yaw rate, resulting in heterogeneous collaborative perception fusion with varying state spaces. In [220], researchers introduced a CI-based track-to-track fusion approach to address tracks with unequal dimensions. This approach involved augmenting the low-dimensional track and exploring various methodologies for state augmentation. The effectiveness of state augmentation significantly impacted the CI’s optimization process, as demonstrated through comparative analysis.
Late collaboration can maintain collaborators’ model information privacy by allowing each agent to train its model individually and share outputs. However, individual outputs may introduce noise, leading to a degradation in cooperative performance. Intermediate collaboration has become a prominent choice for V2X CP due to its flexibility. In real-world scenarios, connected agents may use different feature extractors, resulting in a domain gap among shared features. To mitigate this feature gap, the work in [221] proposed a Multi-agent Perception Domain Adaptation (MPDA) framework. The MPDA framework included a feature resizer and a sparse cross-domain transformer. The feature resizer aligned the spatial and channel dimensions of received intermediate features using the FAX attention block [215] and a convolution layer. The sparse cross-domain transformer generated domain-invariant features through an adversarial strategy to deceive a domain classifier, effectively reducing the feature domain disparities when connected agents employed different feature extraction networks.
Drawing inspiration from the parameter-efficient fine-tuning strategy [248], a Model Adaptation for Cooperative Perception (MACP) was proposed in [222]. Specifically, a single-agent pre-trained model was adapted to the cooperative setting by freezing most of its parameters and adding lightweight modules. Domain shift mitigating was achieved through the design of the Convolution Adapter (ConAda) for the feature encoder and communication channel. The ConAda module helps adapt the pre-trained single-agent parameters to the cooperative setting by projecting the source feature map to a lower dimension and then transforming it back to the source space. This adaptation allows the model to handle spatial distribution shifts and feature space shifts that occur when combining point clouds from different vehicles. The MACP framework effectively addressed domain shifts and enhanced the perception capabilities. However, it may suffer from biases in the pre-trained model, leading to lower average precision for objects in the 50-100m range compared to other methods.
V-A2 Data Heterogeneity
In practical autonomous driving scenarios, connected agents often possess distinct sensory modalities. Agents with LiDAR sensors provide precise and accurate geometrical information, while those equipped with cameras offer comprehensive and contextually rich semantics. Collaborating with agents possessing diverse sensor modalities can enhance the ego’s sensing capabilities by leveraging unique yet complementary environmental attributes captured by each sensor. However, sharing information among camera and LiDAR agents introduces semantic disparities. Additionally, disparities exist between measurements from roadside infrastructures and those from CAVs. Different sensors may exhibit varying noise levels due to factors like maintenance frequency and hardware precision. These distinctive attributes of connected agents present significant challenges in developing a heterogeneous cooperative system.
In a multi-agent hetero-modal setting, connected agents are equipped with various sensors, forming a dynamically heterogeneous graph. The presence and types of these sensors are stochastic, and their relative poses change across different scenes. This necessitates an adaptive architecture that can accommodate diverse agent quantities and typologies while addressing the misalignment of CP data. To address these challenges, HM-ViT [223] facilitated V2V cooperative 3D detection through a Heterogeneous 3D Graph Attention (H3GAT) module. In this module, each node represents a feature vector of an agent’s feature map, and its core function is to extract heterogeneous inter-agent and intra-agent visual cues within the BEV feature space. Node states are updated by aggregating features derived from transformer blocks with a heterogeneous normalization layer, followed by a hetero-modal Multi-layer Perceptron Layer that processes camera and LiDAR features with distinct parameter sets. HM-ViT demonstrated remarkable flexibility and robustness while achieving superior performance in scenarios with varying sensor modalities.
Sensors of identical modalities may exhibit heterogeneity in terms of brand, resolution, and configuration parameters. The overlook of the domain discrepancy engendered by the disparate sensor attributes utilized by each entity can deleteriously impact the culminating efficacy of the system [224]. Consequently, it becomes imperative to devise a mechanism for the extraction of domain-agnostic representations from data originating from multiple sources within the ambit of V2X collaborative perception.
In response to this exigency, a teacher-student distillation architecture, designated as DI-V2X, was introduced in [224]. This framework is characterized by the integration of three principal components: a Domain-Mixing Instance Augmentation (DMA) module, a Domain-Adaptive Fusion (DAF) module, and a Progressive Domain-Invariant Distillation (PDD) module. The DMA module was deployed to mediate the distributional alignment between the teacher model and the student models, thereby laying the groundwork for an efficacious distillation endeavour. Concurrently, the DAF module amalgamates features extracted from both vehicles and infrastructure, engendering a composite representation enriched with pertinent information from each respective domain. In a parallel vein, the PDD module orchestrates the alignment of disparate student features with the teacher’s BEV feature, both before and after fusion, under the directive of overlapping and non-overlapping masks. Through the synergetic operation of these modules, DI-V2X substantiated an enhancement in V2X perception tasks, fostering a generalized representation that reconciles point cloud data across a spectrum of domains.
Besides sensor characteristics, diverse environmental conditions inevitably create a substantial domain gap and data heterogeneity among CAVs [225]. To address these issues, the work in [225] proposed a domain generalization framework comprising an Amplitude Augmentation (AmpAug) method, an intra-system domain alignment mechanism, and a meta-consistency training scheme. The AmpAug is to increase the diversity of low-frequency variations, forcing the model to obtain domain-invariant representations. This enhancement can reduce the model’s influence on low-spectral variations in the image. The meta-consistency training scheme was utilized to simulate domain shifts and optimize the model with consistency loss. The domain alignment mechanism translates image styles and unifies the distribution of image pixels in colour space among CAVs before the inference stage, potentially eliminating the domain discrepancy among CAVs prior to inference.
Furthermore, the distinctness in data characteristics caused by different types of agents also poses significant challenges in designing a heterogeneous cooperative system. VINet [226] addressed this issue using a Two-stream Fusion (TSF) algorithm to integrate heterogeneous features captured from infrastructures and CAVs on an edge server. Specifically, a Lightweight Feature Extraction (LFE) module reduced computing loads in feature extraction without involving the CNN backbone. TSF employed two feature streams with different LFEs to extract pillar features from CAVs and connected infrastructures. Each feature stream implemented a feature mapping and max-out procedure to calculate fused features. The resulting fused features from the two streams were concatenated and interpreted by a CNN backbone to generate detected bounding boxes. VINet proved effective for cooperative object detection with low system-wide costs under dynamic, scalable, and heterogeneous conditions.
V-B Data Rate Requirements
The data rate requirements for CP are expected to increase significantly with the advancement of automated driving levels [249]. Reliable communication is a critical necessity for CDA, as it involves continuously and timely updating dynamic information about surrounding moving obstacles to prevent traffic accidents. Therefore, determining the minimum required sensor data rate for ensuring safe autonomous driving becomes imperative.
Organizations like the 3rd Generation Partnership Project (3GPP) and the 5G Automotive Association (5GAA) have presented data rate requirements to support V2X CP-related applications [250, 251]. However, these considerations did not account for safety concerns such as cyber-attacks and software errors. In a study conducted in [227], the minimum required data rate for a CAV to safely navigate through an intersection was analyzed in scenarios involving both automated and human driving. Safety was assessed by distinguishing cases where braking is necessary to prevent a collision, thereby reducing unnecessary braking while ensuring safe driving and traffic efficiency. The minimum required data rate for each velocity category was determined based on case classification and compared with the data rate achievable through conventional V2I and millimeter-wave communication bands. The findings indicated that transmission using the new millimetre-wave communication standards could meet the minimum data rate requirement for safe autonomous driving in such scenarios.
V-C Lossy Communication
In real-world traffic scenarios, lossy communication is common in V2X communication due to various factors. These factors include Doppler shift induced by fast-moving vehicles [252], dynamic topology leading to routing failures [253], adverse weather conditions, and obstacles causing multi-path effects and signal power fluctuations, resulting in packet loss [254]. Additionally, shared data may face interference from other communication networks or be tampered with by attackers before reaching their intended destination. These factors collectively contribute to incomplete or inaccurate shared information during V2X communication, consequently diminishing the performance of CP systems.
Drawing inspiration from image-denoising networks, the Lossy Communication Aware Repair Network (LCRN) [228] was designed to generate tensor-wise filters and recover damaged features received from cooperative CAVs. The LCRN employs an encoder-decoder architecture with skip connections to produce per-tensor filter kernels, which are then applied to the received lossy features. The recovered intermediate features subsequently undergo feature fusion through a V2V attention module. The LCRN effectively mitigates the impact of artifacts resulting from damaged features by employing tensor-wise filtering, although it does not explicitly address feature loss. In contrast, the V2X-INCOP [229] leverages historical information to recover missing data caused by interruptions in V2X communication. The core of this approach is an adaptive communication model capable of extracting multi-scale and spatial-temporal features from historical data to estimate missing information. Empirical results have demonstrated that V2X-INCOP can effectively alleviate the adverse effects of communication interruptions on CP.
V-D Data Trustworthiness
Advanced onboard sensors in intelligent vehicles typically provide a high level of reliability for the overall perception system, allowing the vehicle to be aware of its surroundings. However, these sensors are not immune to challenging scenarios like adverse weather conditions and equipment malfunctions. Similarly, various agents within the V2X communication network exhibit changing performances and may possess different equipment grades. Consequently, inconsistencies in transmitted information are inevitable during V2X CP processes. Furthermore, malicious agents may tamper with shared packets or generate forged data [255]. Ego vehicles lack the capability to independently verify events they might have missed and ascertain the authenticity of received data [256]. Overreliance on such misleading information by driver assistance systems can disrupt their functioning, posing risks to road users. Therefore, assessing the quality of received information and detecting erroneous data before proceeding with further processing is essential.
To address this concern, probabilistic modelling techniques were employed in [230] to estimate the trustworthiness of a pair of connected agents based on the consistency of the data they provide. The underlying concept involved quantifying the consistency of received data using the Bayes filter, enabling the identification of forged information based on the relationship between data consistency and agent trustworthiness. However, this method can only detect attackers when there is a substantial redundancy (at least triple redundancy) in the contradictory data.
Recognizing that spoofing attacks often target a single vehicle at a time, FDII [231] capitalized on LiDAR data from neighbouring vehicles to counteract spoofing attacks on CAVs. By comparing the victim vehicle’s LiDAR scan with unaltered scans from others, FDII could discern the additional points introduced in the victim’s LiDAR data. The differences in detected obstacles under each scan were analyzed using a decision tree algorithm to classify various attacks and update the unsafe areas for each vehicle. The results demonstrated FDII’s effectiveness in accurately identifying attack types and reconstructing vulnerable areas. However, it is worth noting that decision trees may struggle to capture complex data relationships and tend to favour the majority class in imbalanced datasets, potentially leading to suboptimal classification of minority classes.
To mitigate data fabrication attacks, the Collaborative Anomaly Detection (CAD) model [232] enabled benign vehicles to jointly reveal malicious fabrication by sharing occupancy maps which are then merged into one. The occupancy maps are represented using fine-grained polygons, which provide greater flexibility and precision compared to grid-based occupancy maps. Motion estimation is also performed to track objects over consecutive frames and attach motion information to the occupancy map. The consistency checks are then implemented to reveal inconsistencies among synchronized occupancy maps and compare the perception results with the merged occupancy map. Alerts are raised on conflicted regions, indicating potential anomalies or attacks. Although CAD has been shown to be effective in detecting and mitigating attacks in real-world scenarios, it has certain limitations, such as its effectiveness depending on the presence of at least one benign CAV observing the attacked region.
Malicious Agent Detection (MADE) [233] was designed for multi-agent cooperative perception networks, leveraging two statistics, match loss and collaborative reconstruction loss, to detect malicious agents. The match loss measures the difference in output bounding box proposals between the ego agent and the agent being inspected, while the collaborative reconstruction loss assesses the consistency of feature maps when the ego agent collaborates with the agent being inspected. These statistics can be used independently or jointly to construct a multi-test with a controlled false detection rate using the Benjamini-Hochberg procedure [257]. This study provided valuable insights into the detection of malicious agents. However, the detection inference in MADE is based on a multi-test using conformal p-values computed on calibration sets. This approach relies on the assumption that the ego agent is guaranteed to be benign, which may not always hold true in real-world scenarios.
V-E Perception Uncertainty
Multi-agent cooperation significantly enhances the perception efficiency of autonomous vehicles. However, inherent uncertainties persist in CAVs due to out-of-range objects, sensor errors, and adverse weather conditions [258]. These uncertainties can result in false detections, adversely affecting subsequent self-driving modules like planning and control. Addressing these uncertainties within CP is essential for ensuring safe autonomous driving.
Accurately capturing the uncertainties of the perception system represents a promising approach to enhancing CAV safety. The Perception Error Model (PEM) [259] introduced single-agent perception errors into the ground truth to directly model perception uncertainties within the simulation pipeline, bypassing the need for synthetic sensor data. In an expansion of this concept, the Cooperative Perception Error Model (CoPEM) [234] was introduced to tackle occlusion issues by modelling V2X-relevant perception errors and uncertainties. CoPEM facilitates the tuning of perception performance in simulation and scales effectively with an increasing number of sensors for CP. However, it necessitates a statistical model for perception errors for each connected agent and does not account for the high maneuverability of vehicle targets. To address this limitation, the Consensus-based Distributed Multi-Model CP (DMMCP) [235] approach was proposed. DMMCP enhances the model uncertainty of maneuvering targets by integrating prior knowledge from diverse motion models. It employs a hybrid consensus strategy within a multi-layer information fusion framework, combining it with an advanced Cubature Kalman Filter (CKF) [260] to incorporate more precise nonlinear system models and sensor data. DMMCP demonstrates significant effectiveness, particularly when considering model-related uncertainties. Nevertheless, initializing the CKF can be challenging, especially when dealing with uncertainties in the initial state estimate.
While CoPEM and DMMCP provide probabilistic perception results that offer partial information, the GevBEV [171] presents a comprehensive probabilistic model. GevBEV captures perception uncertainty by constructing a probabilistic BEV map of the 2D driving space using point-based spatial Gaussian distributions. It leverages Evidential Deep Learning (EDL) [184] to estimate probabilities and associated uncertainties for new sample points in the driving space. The model interprets the sampled Gaussian distribution densities as evidentiary support for the Dirichlet distribution. Consequently, GevBEV provides a quantifiable and explainable level of uncertainty for its predictive outcomes. Experimental results indicate that GevBEV outperforms previous research on the benchmark OPV2V for BEV map interpretation.
Uncertainty quantification is critical for safety-critical systems in CAVs, helping determine the likelihood of specific outcomes when certain system aspects are imprecisely known. The Double-M quantification method [236] estimates the uncertainty in cooperative object detection by leveraging temporal features in the dataset. At its core, Double-M employs a customized Moving-Block Bootstrap (MBB) algorithm to model the multivariate Gaussian distribution of each corner of the bounding box. MBB accounts for auto-correlations within time-series data by sampling data from constructed data blocks during training. Double-M quantification assigns significant uncertainties to low-accuracy objects, contributing to later modules in self-driving tasks. While Double-M reduces uncertainty and enhances accuracy, it relies on the assumption that each corner of the bounding box adheres to an independent multivariate Gaussian distribution.
In the context of object tracking, the MOT-CUP framework [237] was designed to address uncertainty quantification for cooperative object detection. MOT-CUP incorporates detection uncertainty into the Kalman Filter and the association procedure for object tracking. It quantifies uncertainty in cooperative object detection through direct modelling [261] and conformal prediction [262]. This obtained uncertainty information is then propagated to the Standard Deviation-based Kalman Filter (SDKF) and the Negative Log-likelihood (NLL) based association improvement procedure for multiple object tracking. SDKF uses the uncertainty information to improve location prediction precision and track state prediction. NLL-based metrics are introduced to consider the uncertainty of collaborative detection while handling low-quality detection results. Another 3D multi-object cooperative tracking algorithm, DMSTrack, that utilizes a differentiable multi-sensor Kalman Filter to estimate measurement uncertainty for each detection, was introduced by [238]. The Kalman Filter prediction step used a constant velocity model as the process model. A covariance neural network was designed to learn the observed noise covariance for each detection from a different CAV by using local point cloud features and positional features as inputs. DMSTrack estimates measurement uncertainty for each detection to better utilize the theoretical optimality property of Kalman Filter-based tracking methods. Extensive experiments demonstrate the potential benefits of incorporating uncertainty quantification from CP into multiple object tracking.
V-F Task Discrepancy
Most existing CP approaches are task-specific, necessitating retraining the entire model for different perception tasks. These task-specific methods have limitations and potentially hinder the models’ generalization ability across a broader spectrum of perception tasks. The development of a CP framework independent of downstream tasks holds promise for advancing cooperative driving automation. In this context, recent endeavours have focused on a novel task known as multi-agent scene completion, also referred to as cooperative reconstruction. In this task, each agent learns to efficiently share information to reconstruct a complete scene observed by all agents. By training models to reconstruct holistic observations, this approach encourages the acquisition of more effective task-agnostic feature representations.
The spatiotemporal auto-encoder (STAR) [239] aims to disregard downstream tasks and concentrate on learning to reconstruct an entire scene through shared features in a self-supervised manner. STAR facilitates the transmission of sub-sampled features to collectively cover the entire spatial region, striking a balance between scene reconstruction performance and communication costs. The reconstructed scene can then be used for various downstream tasks without additional training. This task-agnostic methodology represents a novel direction in CP, enabling the separation of cooperation training from downstream task learning. However, due to its imperfect scene completion, STAR still exhibits a performance gap in downstream perception compared to task-specific methods.
In contrast to STAR’s self-supervised learning approach, the CORE method, as described in [240], learns to reconstruct comprehensive scenes under the guidance of BEV representations derived from the aggregation of raw sensor data from all agents. When provided with a group of agents and their 2D BEV maps, CORE conducts cooperative reconstruction through three main components: a compression module, a collaboration module, and a reconstruction module. The compression module calculates a compressed feature representation for each BEV, considering channel-wise compression and sub-sampling features along the spatial dimension. Drawing inspiration from prior work [172], CORE incorporates a lightweight attention-aware collaboration module to facilitate information aggregation. The reconstruction module employs a decoder structure to regenerate a complete scene observation from the fused features. This “learning-to-reconstruct” approach is task-agnostic and provides clear and meaningful supervision to encourage more effective cooperation, ultimately benefiting perception tasks.
V-G Data Privacy
Despite the promising outcomes of collaboration among multiple agents, it necessitates local data exchange among cooperating partners. In real-world scenarios, connected agents often exhibit heterogeneity, and privacy concerns arise as automotive manufacturers and software providers are typically reluctant to share private data. This reluctance results in incomplete and insufficient information for perception tasks. Addressing privacy issues in collaborative driving automation is thus of paramount practical significance.
Federated Learning (FL)-assisted CAVs represent an emerging paradigm aimed at mitigating privacy concerns while reducing communication costs and diversifying the seen scenarios in comparison to conventional centralized learning [263, 264]. FL enables collaborating clients to effectively train models without the need for data aggregation or direct data exchange [265]. However, determining the network resources and sensor placements for multi-stage training with multi-modal datasets in varied scenarios can be a complex challenge.
Conventional FL approaches have limitations in handling data heterogeneity stemming from diverse sensor configurations. Variations in camera positions across different clients can lead to FL results that deviate from the optimal values of local data. To address this challenge, the FedBEVT framework [242] introduces a federated transformer learning approach grounded in camera-attentive personalization. This approach mitigates the adverse effects of data heterogeneity by privatizing the camera positional embedding. Empirical investigations have demonstrated the efficacy of FL techniques in training a transformer-based model to enhance BEV perception. Nevertheless, FedBEVT does not account for network heterogeneity. Varied connection qualities among clients can lead to communication delays during global model aggregation, hindering the FL process.
To tackle the complexities of data and network heterogeneity in vehicular networks, the work in [243] proposed a contextual client selection pipeline. This pipeline leverages digital twin simulations to emulate networks and calculates FL communication latency based on predictive transport scenarios. Specifically, it groups CAVs with similar data distribution and employs the Fast-$\gamma$ rule to select a subset of clients ($\gamma$ clients) with the lowest communication delays per cluster. This FL approach effectively addresses both data and network heterogeneity, enhancing communication efficiency through contextual client selection.
Challenges in FL for CAVs also include unfairness due to differences in data volumes and communication delays. An FL algorithm to address challenges related to imbalanced data distribution and fluctuating channel conditions in CAVs was introduced in the work [241]. A local training strategy was designed specifically for each CAV based on their data volume and channel conditions to reduce the number of global rounds required and shorten the time intervals between each round. This allows it to achieve fairness in energy and time costs, accelerating model convergence while maintaining a balanced cost for all CAVs.
V-H Simulation to Reality
Due to the scarcity of real multi-agent data and the resource-intensive nature of manual labelling, most existing CP models often rely on simulated sensor data for training and testing. However, when these simulation-trained models are deployed in real-world settings, they frequently experience a decline in perception performance due to the substantial domain gap between simulation and reality. Domain adaptation techniques are employed to address these challenges by adapting models trained on labelled source domains to perform effectively in unlabeled target domains.
In contrast to the idealized simulation environment, real-world multi-agent scenarios introduce localization errors and communication latency among agents. Furthermore, the feature distribution in actual driving situations is notably more intricate than that in simulated data. To mitigate these two prominent disparities, the work in [244] introduced the Simulation-to-Reality (S2R) transfer learning framework, specifically S2R-ViT, which leverages labelled simulated data and unlabeled real-world data to minimize domain discrepancies in cooperative 3D object detection. The S2R-ViT framework comprises two key components: the S2R Uncertainty-aware Vision Transformer (S2R-UViT) and the S2R Agent-based Feature Adaptation (S2R-AFA). In S2R-UViT, two multi-head self-attention branches (LG-MSA) are utilized to incorporate both local and global attention mechanisms, enhancing feature interactions across spatial positions of all agents. Following LG-MSA, an uncertainty-aware module is introduced to assign weights to the ego-agent’s features while considering uncertain factors stemming from the shared features of other agents. The S2R-AFA incorporates inter-agent and ego-agent discriminators, enabling S2R-ViT to generate robust, domain-invariant feature representations. Experimental results underscore the importance of domain adaptation for LiDAR-based 3D object detection when transitioning from simulation to real-world environments.
In addition to the domain gap between simulated and real-world data, there is also a domain gap among real-world collaborative agents, which further complicates the challenge of sim2real generalization. To tackle these issues, an unsupervised sim2real domain adaptation method named Decoupled Unsupervised Sim2Real Adaptation (DUSA) was proposed in [245]. The DUSA solves the problems of sim2real adaptation and inter-agent adaptation by introducing the Location-adaptive Sim2Real Adapter (LSA) module and the Confidence-aware Inter-agent Adapter (CIA) module, respectively. The LSA module uses location importance to prompt the feature extractor to produce sim/real invariant features, while the CIA module uses confidence cues to help the feature extractor output agent-invariant features. This method effectively bridges the gap between simulated and real-world data for V2X collaborative perception, demonstrating its superiority over existing methods through experiments.
VI Issues for Dataset
While CP has gained momentum in recent years, it remains a nascent field. One of the key obstacles impeding the progress of CP algorithms is the substantial cost and potential safety risks associated with conducting field experiments. Such experiments necessitate deploying numerous expensive CAVs and sizeable test areas. Additionally, acquiring precise annotations requires significant labour and time, especially when dealing with CP involving multiple agents [266]. To address these challenges efficiently, an effective research strategy is to develop and validate CP algorithms within a simulated environment, which offers a cost-effective and safe alternative to real-world experimentation. However, validating V2X perception in authentic scenarios remains challenging due to the absence of well-established public benchmarks. This section delineates prominent simulation platforms and open-source datasets frequently employed in developing V2X CP systems.
VI-A Simulation Tools
Numerous simulation platforms have been developed to offer multi-modal sensory inputs and 3D annotations, thereby advancing the development of multi-agent CP before the widespread availability of real-world datasets. For example, SUMO [267] is a traditional traffic and driver behaviour simulator renowned for its capacity to handle large-scale and realistic traffic flows. It incorporates dynamic modelling for each vehicle, enabling users to swiftly create customized traffic scenarios through the Traffic Control Interface (TraCI) API. SUMO is highly convenient for integrating naturalistic trajectory data, which can be directly applied to the surrounding environment for testing CDA. However, it has limitations when providing high-resolution sensor data with exceptional fidelity.
In recent years, there has been a proliferation of autonomous driving simulators designed to facilitate high-fidelity modelling of the traffic environment and sensor capabilities. The following content will provide an overview of the fundamental characteristics of several prominent simulation tools used to develop and analyze CDA systems.
VI-A1 CARLA
CARLA [268] is a pioneering autonomous driving simulator celebrated for its versatility in specifying sensor suites and environmental conditions. This open-source platform was purpose-built to facilitate development, training, and validation of novel autonomous driving systems. CARLA’s architecture is notably scalable, employing a server-multi-client model to distribute computations across multiple nodes. The server continuously updates the environmental physics while users exert control over the client side through the CARLA API. Although CARLA permits environment extension, the process of swiftly creating scenarios and managing substantial traffic volumes has grown intricate. CARLA does offer a traffic manager module for generating background traffic. However, these modules rely on simplified behavioural rules that may not faithfully replicate actual driver behaviour.
VI-A2 OpenCDA
OpenCDA [269, 270] stands as an open and dedicated co-simulation framework designed for CDA. It encompasses both traffic and vehicles, offering a complete CDA vehicle software pipeline. OpenCDA seamlessly integrates CARLA and SUMO for lifelike scene rendering, vehicle modelling, and traffic simulation. This framework facilitates the evaluation of CDA within a CARLA + SUMO co-simulation environment and supports diverse levels and categories of cooperation among CAVs during simulation. OpenCDA also provides a comprehensive library of CDA research pipelines, source code, and standard autonomous driving components encompassing perception, localization, planning, control, and V2X communication modules. One of its key advantages is its flexibility, allowing users to easily substitute default modules with custom-designed ones to assess the impact of new functionalities on overall CDA performance.
VI-A3 V2XP-ASG
V2XP-ASG [271] is designed to assess LiDAR-based V2X perception models by providing the capability to generate complex and realistic scenarios. This simulation framework comprises two key components: Adversarial Collaborator Search (ACS) and Adversarial Perturbation Search (APS). The ACS creates an adversarial collaboration graph based on an existing scene. It identifies and selects collaborators whose combined perspectives will likely lead to sub-optimal outcomes in perception tasks. The APS, on the other hand, focuses on perturbing the positions of vehicles within the scene. It achieves this using black-box optimization algorithms tailored to perception tasks’ adversarial objectives. Combining these two stages in V2XP-ASG offers an efficient approach to enhance the complexity and difficulty of the scenarios presented in the simulation, thereby challenging the V2X perception models under evaluation.
VI-A4 AUTOCASTSIM
In [165], researchers introduced AUTOCASTSIM, an autonomous driving simulation framework built on the CARLA platform. AUTOCASTSIM was designed to evaluate cooperative driving models, particularly in accident-prone scenarios. This platform offers three benchmark traffic scenarios (Overtaking, Left Turn, and Red Light Violation) that impose decision-making challenges due to restricted line-of-sight sensing. AUTOCASTSIM includes an integrated networking simulation feature, facilitating customization of V2V communication dynamics. Additionally, it incorporates an advanced driving model that incorporates privileged environmental information, enabling the creation of customized and realistic traffic scenarios.
VI-A5 CMM Co-simulation Platform
The work in [272] highlighted a common limitation in most current traffic simulators, which typically use intrinsic attributes of target objects without considering imperfect perception in CDA models. A CARLA-based co-simulation platform called the Cyber Mobility Mirror (CMM) was developed to address this limitation. The CMM platform enhances CDA by generating realistic perception data. A deep learning-based perception pipeline was integrated into the simulation system to establish a CMM framework capable of faithfully perceiving and reconstructing simulated traffic objects in 3D. This improvement enables CDA models to utilize perceived information from the CMM system, enhancing model fidelity and supporting more robust system validation, ultimately boosting overall confidence.
TABLE VIII: Comparison between Cooperative Perception-related Datasets.
Dataset Source Scenario Sensors Tasks
Camera LiDAR Depth IMU/GPS Object Detection Object Tracking Segmentation Motion Prediction Accident Prediction
DAIR-V2X-C [273] Real-world V2I $\surd$ $\surd$ $\times$ $\surd$ $\surd$ $\times$ $\times$ $\times$ $\times$
V2X-Seq [274] Real-world V2I $\surd$ $\surd$ $\times$ $\surd$ $\surd$ $\surd$ $\times$ $\surd$ $\times$
V2V4Real [275] Real-world V2V $\surd$ $\surd$ $\times$ $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\times$
LUCOOP [276] Real-world V2V $\times$ $\surd$ $\times$ $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\times$
TUMTrafV2X [277] Real-world V2I $\surd$ $\surd$ $\times$ $\surd$ $\surd$ $\surd$ $\times$ $\surd$ $\surd$
HoloVIC [278] Real-world V2X $\surd$ $\surd$ $\times$ $\surd$ $\surd$ $\surd$ $\times$ $\surd$ $\times$
CODD [279] CARLA V2V $\times$ $\surd$ $\times$ $\times$ $\surd$ $\surd$ $\times$ $\times$ $\times$
OPV2V [280] CARLA & OpenCDA V2V $\surd$ $\surd$ $\times$ $\surd$ $\surd$ $\surd$ $\times$ $\times$ $\times$
V2X-Sim [281] CARLA & SUMO V2X $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\times$ $\times$
V2XSet [198] CARLA & OpenCDA V2X $\surd$ $\surd$ $\times$ $\times$ $\surd$ $\surd$ $\times$ $\times$ $\times$
DOLPHINS [282] CARLA V2X $\surd$ $\surd$ $\times$ $\times$ $\surd$ $\times$ $\times$ $\times$ $\times$
DeepAccident [283] CARLA V2X $\surd$ $\surd$ $\times$ $\times$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
VI-A6 OpenCDA-ROS
Existing CDA simulation platforms often exhibit disparities between simulation and real-world scenarios. To address this issue comprehensively, a unified tool and framework are needed to support CDA development in simulation and real-world environments. OpenCDA-ROS [284] has been developed as an extension of OpenCDA, tailored for the Robot Operating System (ROS). This adaptation serves as a bridge between simulated and real-world CDA, ensuring that algorithms and vehicles perform similarly in both settings. By reorganizing cooperative functions into a more shareable structure, OpenCDA-ROS benefits simulation and real-world prototyping and deployment, facilitating the seamless integration of these two domains in CDA development.
VI-B Open Source Dataset
The development of CP models relies on perception datasets gathered from various onboard sensors like radar, camera, and LiDAR. Utilizing these sensors for real-world perception tasks can be costly and time-consuming. A more economically viable alternative is to obtain cooperative data from simulation platforms featuring high-fidelity sensor modelling and perception capabilities. This subsection will outline several datasets relevant to CDA research, summarized in Table VIII.
A direct approach to evaluating V2X CP involves collecting diverse real-world test scenarios. DAIR-V2X-C [273] is an autonomous driving dataset from real scenarios designed for research on V2I cooperation-related challenges. It comprises 71,254 LiDAR and 71,254 Camera frames captured at intersections equipped with infrastructure sensors. Expert annotators have meticulously labelled all frames with 3D annotations. While DAIR-V2X supports the perception task, it does not cover the essential motion prediction aspect.
V2X-Seq [274] offers a sequential V2I dataset, also collected from real-world scenarios, serving to develop and validate CP and motion forecasting algorithms. This dataset comprises two distinct sub-datasets: the sequential perception and trajectory forecasting datasets. The former includes over 15,000 frames from 95 scenarios, encompassing various data types such as infrastructure images, point clouds, vehicle-side images, point clouds, 3D detection/tracking annotations, and vector maps. The latter contains 210,000 scenarios, including 50,000 cooperative cases from 28 intersections, with a substantial data coverage of 672 hours.
For V2V scenarios, the V2V4Real dataset [275] was collected using two vehicles equipped with multi-modal sensors, driving through diverse scenarios in Columbus, Ohio, USA. In comparison to DAIR-V2X and V2X-Seq, V2V4Real covers a more extensive driving area and mileage, encompassing four road types: intersections, highway entrance ramps, highway straight roads, and city straight roads. This dataset comprises 40,000 RGB frames, 20,000 LiDAR frames, 240,000 annotated 3D bounding boxes for five object classes, and corresponding HDMaps for road topology prediction and semantic understanding of the BEV map. This substantial real-world dataset facilitates the training and testing of CP models for cooperative 3D object detection and tracking, as well as Sim2Real domain adaptation. It also offers comprehensive benchmarks involving eight state-of-the-art CP algorithms for these three tasks.
The V2V4Real dataset, with only two vehicles, lacks comprehensiveness for advancing research on cooperative applications. Addressing this limitation, LUCOOP [276] presents time-synchronized multi-modal data collected by three CAVs, each equipped with a broad sensor suite. This suite includes at least one LiDAR sensor, one GNSS antenna, and up to three IMUs. Additionally, Ultra-Wide-Band (UWB) sensors are mounted on each vehicle and placed statically along the trajectory to enable both V2V and V2X range measurements. The dataset encompasses sensor measurements in a multi-vehicle setup, spanning a trajectory of over 4 km and a time interval of more than 26 minutes. Overall, it includes more than 54,000 LiDAR frames, approximately 700,000 IMU measurements, over 2.5 hours of 10 Hz GNSS raw measurements, and 1 Hz data from a reference station [276]. The dataset also offers more than 6,000 total station measurements over a trajectory exceeding 1 km, as well as 1,874 V2V and 267 V2X UWB measurements [276]. Additionally, it provides 3D bounding box annotations for evaluating object detection methods and highly accurate ground truth poses for each vehicle throughout the measurement campaign [276].
The current practical utility of the LUCOOP dataset is limited by the absence of complex traffic scenarios. In contrast, the TUMTraf-V2X dataset [277] presents a comprehensive resource comprising 2,000 annotated point clouds and 5,000 annotated images collected from a variety of sensors. This dataset encompasses a total of 30,000 3D bounding boxes with associated track IDs, with a particular emphasis on challenging scenarios like near-misses, overtaking, U-turns, and traffic violations. Data from nine distinct sensors are included, capturing traffic scenes under diverse weather and lighting conditions. Additionally, the dataset includes rare occurrences, such as pedestrians crossing busy intersections against red lights. An open-source annotation tool and development kit are provided to facilitate the integration and evaluation of the dataset.
To address the absence of extensive datasets for real-world V2X communication, the study by [278] introduced a multi-modal dataset of ETSI ITS V2X messages gathered from public road traffic. This dataset includes over 230,000 ETSI ITS V2X messages from a diverse set of more than 1,800 vehicles and roadside units. The study provides details on the hardware setup and recording locations used to collect the V2AIX dataset, which encompasses urban and highway traffic sections in Germany. Furthermore, it provides insights into the current status of standardized V2X communication and offers a way to integrate these messages into the Robot Operating System (ROS) for automated driving applications.
While realistic datasets allow CP models to generalize effectively to real-world driving situations, they are expensive, time-consuming, and limited in scenarios. A cost-effective alternative is to gather large-scale datasets in a high-fidelity simulator. The official version of the Cooperative Driving Dataset (CODD) [279], generated using the CARLA simulator, consists of 108 snippets. Each snippet contains 125 temporal frames of LiDAR point cloud data. Each frame in a snippet corresponds to a temporal data slice, including LiDAR data from all vehicles in the scenario, sensor absolute pose information, and ground-truth annotations for 3D bounding boxes of cars and pedestrians. CODD focuses on a single driving scenario. In contrast, the OPV2V dataset [280] incorporates camera images and LiDAR point clouds from connected vehicles across various scenes. Using OpenCDA and the CARLA simulator, OPV2V collects over 70 diverse scenes featuring varying numbers of connected vehicles. These scenes span eight towns within CARLA and a digital town, Culver City, Los Angeles. CODD and OPV2V provide ground truth for cooperative 3D object detection with V2V communication. However, neither dataset accounts for roadside infrastructure, limiting their ability to evaluate V2I perception.
The V2X-Sim dataset [281] serves as a synthetic V2X-assisted CP dataset in the field of autonomous driving. It is designed to stimulate research in multi-agent, multi-modality, and multi-task perception. The dataset was generated by SUMO to create numerically realistic traffic flows and CARLA to capture sensor streams from multiple vehicles at the same intersection. V2X-Sim offers well-synchronized recordings from roadside infrastructure and multiple vehicles at the intersection, along with comprehensive ground truth annotations to support various downstream tasks, including detection, tracking, and segmentation. However, it lacks noise simulation in its settings and only encompasses a single category of the road network.
In contrast, the V2XSet dataset [198] explicitly accounts for real-world noise during V2X collaboration by leveraging CARLA and OpenCDA. It includes more than 11,447 frames divided into three subsets: training, validation, and testing, with ratios of 6,694, 1,920, and 2,833, respectively. The DOLPHINS dataset [282], generated using the CARLA traffic simulator, encompasses six distinct autonomous driving scenarios. It supports V2X communication and provides temporally synchronized sensor data from the ego vehicle, a cooperating vehicle, and a roadside unit. The dataset comprises 42,376 frames of sensor data and 292,549 3D bounding box annotations for cars and pedestrians.
While the datasets above cover large-scale scenarios, they lack accident scenarios essential for comprehensive safety assessments of autonomous driving. Addressing this need, the DeepAccident dataset [283] was created through the CARLA simulator. The DeepAccident dataset comprises 57,000 frames and 285,000 samples, enabling the evaluation of motion and accident prediction capabilities. In each scenario, four vehicles and one infrastructure unit are established to capture sensor data, including multi-view RGB cameras and LiDAR. The dataset includes corresponding annotations for perception and motion prediction tasks. This dataset fills the gap by providing safety-critical scenarios absent in existing V2X datasets, offering diverse viewpoints for accident scenarios. Additionally, an end-to-end accident prediction task is introduced to forecast collision accidents, including their occurrence, timing, location, and the involved vehicles or pedestrians. The proposed V2XFormer serves as a benchmark for motion and accident prediction, as well as object detection, demonstrating superior performance in these predictive tasks compared to single-vehicle models.
VII Lessons Learned
CP technology plays a pivotal role in overcoming the limitations inherent in single autonomous vehicles, addressing issues like restricted perception range and obscured areas. However, the unique characteristics of autonomous driving and the evolution of communication technology bring forth a multitude of challenges for CP. Drawing from recent research and comprehensive analysis of cooperative autonomous driving, we have gleaned valuable insights and identified key open challenges that will significantly influence the future research trajectory in this field. In this section, we will delve into the foremost among these insights, providing an in-depth examination of the hurdles that V2X CP must overcome to fully realize its anticipated potential.
VII-A Open Challenges
VII-A1 Collaboration Trigger
Much of the existing literature predominantly focuses on delineating cooperative behaviours among agents within specific scenarios, often neglecting the essential aspect of empowering the ego CAVs to discern opportune moments and suitable contexts for collaboration. To facilitate this, CAVs require a collaboration trigger mechanism capable of ascertaining the potential benefits of solo travel versus cooperative interaction with other agents at a given time step and within the prevailing driving conditions. The construction of effective collaboration triggers requires exploring pertinent decision factors, including but not limited to weather conditions, road terrain characteristics, and the state of onboard sensors. Furthermore, it is imperative to consider the temporal aspect of cooperation, ensuring that its duration aligns with the preservation of string stability among multiple agents. A compelling avenue for further research involves investigating whether the advantages of cooperation should be anticipated and communicated to CAVs in advance and whether this proactive knowledge dissemination would influence the overall performance of cooperative endeavours. In light of these research gaps, future endeavours are encouraged to tackle the complexities inherent in cooperative decision-making processes.
VII-A2 Cooperative Perception Under Constraints of Realistic Communication
Existing methods for CP have traditionally focused on optimizing the perception capabilities, often disregarding or simplifying the critical constraints imposed by communication factors in the real world. Yet, the real-world communication constraints can heavily degrade the CP performance. Thus, such constraints have to be mimicked realistically. For instance, the limited communication bandwidth emerges as a pivotal constraint, significantly impacting real-time collaboration by increasing transmission time in practical scenarios or reducing the number of bits that can be shared over wireless link. Furthermore, shared data is susceptible to interference from other agents during transmission, exacerbated by obstacles between transmitters and receivers, causing multi-path effects and fluctuation of signal power [285]. Selecting agents with sub-optimal communication interference and poor-quality wireless propagation channels can result in a diminished Signal-to-Interference plus Noise Ratio, reduced data rates, and elevated Bit Error Rates. Consequently, the information transmitted for CP may be discarded or incorrectly decoded. In light of these challenges, it becomes imperative to devise CP technologies that holistically integrate realistic communication constraints modelled by proper networking simulators, detailed-oriented cross-layer communication system models and more reliable wireless signal propagation models [286, 287], alongside the goal of maximizing perception capabilities in practical applications.
VII-A3 Malicious and Selfish Behavior
Reliable communication is of utmost importance to foster effective collaborative driving. Nevertheless, it remains plausible that some agents may occasionally engage in self-serving behaviour, collaborating solely to reduce their costs while causing detriment to others. These self-interested individuals, commonly referred to as “free-riders”, exploit the information shared by other agents to reap the benefits of coordinated efforts without contributing their insights and observations to the collective knowledge pool. Moreover, the distributed deployment of multiple agents diminishes defences against cyber-attacks, rendering them susceptible to malicious activities initiated by untrustworthy users [288]. A malicious collaborator may intentionally transmit erroneous environmental data to neighbouring agents, purposefully dispatch excessive computational tasks, and exhaust valuable computing resources and network bandwidth. The actions of malicious and self-interested agents pose a significant threat to the integrity of the vehicle communication network. The resulting degradation in network services not only undermines communication reliability but also jeopardizes the safety of autonomous vehicles. To fortify the security of the system network, it is imperative for future research to place a heightened emphasis on trust-related concerns among agents. This entails the identification and mitigation of malicious and self-serving behaviours by incorporating them into CP schemes, ideally through the application of suitable methodologies like game theory [289].
VII-A4 Generalizability in Real Scenarios
Constructing real-world scenes and collecting data for researching CP techniques presents substantial challenges. Consequently, many CP methods rely on simulated datasets for development. Although a few recent real-world datasets have become available [273, 275, 274], they are often limited in the number of participating agents, failing to capture the complexities of real-world collaboration scenarios. Therefore, CP frameworks trained on datasets generated through simulation tools may lack optimal generalizability in realistic collaboration settings. Regrettably, the outlook for collecting and annotating perception data in genuine V2X scenarios remains uncertain in the near term. As a viable alternative, it is essential to explore the development of feasible domain adaptation methods [290]. These methods would enable the transfer of CP models between authentic and simulated domains, ultimately facilitating the practical application of V2X CP in various scenarios.
VII-A5 Challenging Scenes and Corner Cases
Ensuring the safety of V2X autonomous driving hinges on the ability to attain robust and precise perception results across diverse and complex scenarios. While several large-scale CP datasets have emerged in recent years, they predominantly concentrate on typical traffic scenarios, often overlooking challenging scenes, notably corner cases. Corner cases encompass atypical and extreme situations within autonomous driving scenarios, marked by uncommon conditions, hardware impediments, and irregular obstacles such as overturned accident vehicles or uniquely shaped vehicles. Due to their scarcity, these corner cases are frequently omitted from perception model training for autonomous driving. Consequently, these models struggle to accurately identify objects, resulting in instances of either missed or erroneous detections, ultimately contributing to severe traffic accidents. To address this pressing issue, there is an immediate need to introduce CP datasets that encompass intricate scenarios and corner cases, bolstering the development of a more resilient CP system.
VII-B Future Direction
VII-B1 Hybrid Collaboration Scheme
Efficient collaboration in multi-agent systems is crucial for CDAs. Early collaboration preserves vital information but requires high-speed communication due to raw data size. Late collaboration with object-level data reduces data size but may degrade performance by discarding valuable information during object proposal merging. Intermediate collaboration, focusing on feature-level data transmission and integration, balances accuracy and bandwidth consumption. However, persistent bandwidth saturation or underutilization challenges exist.
One promising solution involves designing an efficient collaboration trigger. This trigger determines when to cooperate with other agents and selects the most suitable collaborative scheme. Driven by intelligent algorithms and adaptive decision-making, it optimizes bandwidth and resource utilization. Additionally, the hybrid collaboration technique, integrating three fusion schemes, enables agents to partition input frames into raw, feature, and object data segments, facilitating exchange with collaborators. This approach capitalizes on diverse fusion techniques’ strengths and adapts to variable network conditions and environmental contexts. Combining intelligent collaboration triggers with the hybrid collaboration technique holds transformative potential, enhancing multi-agent system efficiency and performance in complex, dynamic scenarios with improved precision and resource management.
VII-B2 Multi-modal Data Sharing
LiDAR sensors offer precise geometric data, while cameras provide rich semantic information. Therefore, a multi-modal perception system could theoretically improve performance by leveraging complementary sensor data. In cooperative autonomous driving, the focus has been mainly on collaboration among agents equipped with identical sensors. However, using different sensor types across various agents can reduce costs while maintaining desirable performance. Integrating multi-modal data from multiple agents presents systematic challenges due to dynamic geometric relationships between LiDAR and camera inputs. Research into CP involving the sharing of multi-modal data is valuable. It is important to note that not all agents possess a suite of multi-modal sensors. Hence, careful consideration of sensor modality is essential when selecting collaborative agents. For instance, prioritizing agents capable of providing high-resolution data for collaboration is advisable during adverse weather conditions.
VII-B3 Integrated Sensing and Communication for Cooperative Perception
Integrated Sensing and Communication (ISAC) is an emerging technology that utilizes the same frequency band and hardware for both sensing and communication [291]. This innovation has the potential to significantly improve spectrum efficiency and reduce hardware costs. In the current context of cooperative autonomous driving, sensing and communication systems typically operate on separate frequency bands. The widespread adoption of millimetre-wave and massive MIMO technologies in infrastructure opens opportunities to implement radar perception in cooperative systems through ISAC. This enhances wireless communication throughput and radar perception resolution in future V2X networks. ISAC is anticipated to be a critical technology in future V2X CP, enabling low-latency, high-data-rate communication, and high-resolution obstacle detection, particularly in challenging scenarios like high occlusion and adverse weather.
ISAC technology can help overcome challenges associated with existing cooperative agent selection methods. It enables dynamic bandwidth allocation and adaptable communication strategies, mitigating limitations imposed by bandwidth constraints, interference, and multi-path effects. This adaptive optimization process ensures efficient and reliable data transmission to its intended destination. Furthermore, ISAC can incorporate error correction and data integrity mechanisms to minimize the risk of data loss or incorrect decoding during transmission. Therefore, ISAC aligns with the need for comprehensive agent selection schemes considering both sensing and communication aspects. It aims to optimize perceptual capabilities while addressing practical challenges related to limited communication resources and potential interference. Ultimately, this advancement enhances the effectiveness of collaborative perception in complex environments.
VII-B4 Responsible AI for Cooperative Perception
Cooperative autonomous driving relies heavily on Artificial Intelligence (AI), which excels in mastering increasingly complex computational tasks. However, AI’s use can also raise concerns, including privacy breaches, bias, and opacity. Balancing AI’s promises with its risks necessitates the adoption of responsible AI principles. These principles, such as fairness, transparency, and privacy, provide guidelines for implementing AI models. While there is a growing focus on responsible AI frameworks, their practical application and impact in V2X CP design have received limited attention. The translation of high-level principles into concrete AI design practices within CP remains unclear.
Responsible AI initiatives can offer valuable insights for creating realistic simulation environments that closely resemble real-world scenarios, benefiting the training and validation of V2X CP frameworks. Additionally, these initiatives can drive the development of domain adaptation techniques, enhancing knowledge transfer between simulated and real domains. This knowledge transfer improves the generalisation of the CP model. Moreover, responsible AI efforts can stimulate the advancement of deep learning methods tailored for generating CP data featuring intricate scenarios and corner cases. Generative models and ethically guided data augmentation techniques are pivotal for constructing comprehensive datasets. This enriched dataset forms the basis for robust training for the V2X CP system, equipping them to navigate and respond effectively to challenging real-world situations.
VII-B5 Privacy-preserved Cooperative Perception
To establish sustainable CP, ensuring the privacy of data and messages exchanged via V2X communications is of paramount importance [292]. Traditional CP systems often rely on centralized servers like RSUs or master vehicles for data fusion algorithms. However, data sharing over V2X communications poses potential risks, as sensitive user information may become exposed to external threats or unauthorized access by servers. FL offers a viable solution to enhance CP intelligence while preserving privacy. FL facilitates intelligent CP by enabling fusion algorithms to run on servers without requiring access to raw data or information sharing. It employs a distributed training mechanism in which collaborating vehicles train localized CP models using their data and then share only the trained model with the server for global aggregation [293]. This approach ensures secure distributed learning, safeguarding the shared model from external adversaries. By prioritizing privacy preservation, FL encourages more CAVs to participate in data perception, thus enhancing the overall robustness of CP systems.
VII-B6 Collaborative AI Integration for Enhanced CP and V2X Communication
In future intelligent transportation systems, efficient collaboration is envisaged between AI systems for V2X communication and AI systems for CP within each CAV.
The integration of AI into radio resource management for vehicle communication systems is driven by the imperative to allocate communication resources efficiently, enabling real-time data exchange among vehicles, infrastructure, and relevant entities [294]. International standardization efforts, led by organizations like the 3GPP, are underway in anticipation of the 6G era. The forthcoming 6G technologies promise unprecedented capabilities in data rates, low latency, and accommodating a massive number of connected devices, all of which are crucial for CP in automated and connected transportation systems.
In the context of CP, each CAV must intelligently determine the information it requires from other connected vehicles and decide what information it should transmit to meet requests from other vehicles. This decision-making process is highly contingent on the dynamic traffic conditions. Therefore, each CAV’s AI system must efficiently manage CP’s necessary information, adapting to the current and near-future traffic environment. However, to optimize information flow management, each CAV must collectively consider the available communication resources at their disposal.
Consequently, further research is imperative to orchestrate effective collaboration and coordination among distinct AI systems operating within individual CAVs and AI systems operating within the V2X network infrastructure. These AI systems must align their objectives and strategies to achieve efficient CP, particularly in dynamic and complex traffic scenarios. Such endeavours are essential to harness the full potential of AI-driven CP for enhancing road safety and traffic efficiency.
VII-B7 Migration of Advanced Perception and Autonomous Driving Framework from Single Vehicle to Cooperative Vehicles
As the development of transformer model [295] in both computer vision and large language model areas, several advanced technologies have been proposed for single-vehicle perception and autonomous driving, e.g. end-to-end perception [296], occupancy grid prediction [297], referring expression-based perception [298], self-supervised world-models for autonomous driving [299, 300], to boost the downstream tasks by using the features extracted from understanding and predicting/reconstruction of the environment, and end-to-end autonomous driving [301], etc. All of aforementioned technologies have great potentials to be extended into cooperative vehicles setting as well, with the expectation for enhancing the capability of autonomous driving. The most recent example is end-to-end autonomous driving under V2X cooperation [302], in which the end-to-end autonomous driving framework is integrated with V2X cooperation to achieve significant performance improvement on planning and other intermediate outputs.
VIII Conclusions
In this paper, we thoroughly examined V2X-based CP and its recent technological progress. Our first focus was on the advancements in V2X communication technologies that support the CP in CAVs. We specifically reviewed recent developments in edge computing, blockchain, digital twin, and 6G that provide low latency, high data rate, and reliable and secure information transmission for CP. Then, we presented a comprehensive CP framework that encapsulates these systems’ fundamental components and core aspects. By adopting a new taxonomic classification for mainstream architectures and conducting in-depth reviews of the literature, we dissected the intricate cooperative challenges that lie at the heart of CP. Furthermore, we highlighted the availability of open-source databases and simulators, which are invaluable tools for evaluating and validating the performance of CP models. These resources facilitate rigorous testing and refinement of CP algorithms in a controlled environment. Moreover, we emphasized the importance of incorporating ethical principles such as fairness, transparency, and privacy into designing and implementing AI models for CP. These principles ensure the ethical deployment of AI in cooperative driving and contribute to the creation of highly realistic simulation environments, domain adaptation methodologies, and enriched datasets that can empower CP models to navigate the complexities of real-world scenarios effectively. In summary, CP is a rapidly evolving area that holds great promise. By combining state-of-the-art V2X communication technology, multi-modal sensing technology, and AI technology for data processing, reliable and efficient CP can be achieved, thus creating safer, more efficient, and more capable autonomous driving and intelligent transportation systems.